### v1

# Aligning Large Language Models with Human Preferences through Representation Engineering

Abstract
Aligning large language models (LLMs) with human preferences is crucial for enhancing their utility in terms of helpfulness, truthfulness, safety, harmlessness, and interestingness. Existing methods for achieving this alignment often involves employing reinforcement learning from human feedback (RLHF) to fine-tune LLMs based on human labels assessing the relative quality of model responses. Nevertheless, RLHF is susceptible to instability during fine-tuning and presents challenges in implementation. Drawing inspiration from the emerging field of representation engineering (RepE), this study aims to identify relevant representations for high-level human preferences embedded in patterns of activity within an LLM, and achieve precise control of model behavior by transforming its representations. This novel approach, denoted as Representation Alignment from Human Feedback (RAHF), proves to be effective, computationally efficient, and easy to implement. Extensive experiments demonstrate the efficacy of RAHF in not only capturing but also manipulating representations to align with a broad spectrum of human preferences or values, rather than being confined to a singular concept or function (e.g. honesty or bias). RAHF’s versatility in accommodating diverse human preferences shows its potential for advancing LLM performance.
1 Introduction
While large language models (LLMs) learn broad-ranging world knowledge and a degree of reasoning proficiency, precise control over their behavior proves challenging due to the unsupervised nature of their pre-training Radford et al. (2018, 2019); Brown et al. (2020); Bubeck et al. (2023); Touvron et al. (2023). For each query, instruction-tuned LLMs Wei et al. (2021); Chung et al. (2022); Touvron et al. (2023) exhibit the capacity to generate multiple responses that are both semantically and syntactically coherent by some sampling techniques. While such ability enables the models to provide diversity that is essential for chat agents, some responses may contain harmful, unethical, socially biased and negative, even illegal content Srivastava et al. (2022); Thoppilan et al. (2022); Bubeck et al. (2023); Wang et al. (2023). Although many responses remain within acceptable bounds, preferences may exist among them, leading to a desire to prioritize specific outputs over others.
Figure 1: Illustration of different apporaches. (a) Reinforcement learning from human feedback (RLHF); (b) Direct preference optimization (DPO); (c) Hindsight instruction relabeling (HIR); (d) Representation alignment from human feedback (RAHF).
Existing methods steer LLMs to align with human preferences often using reinforcement learning (RL), with reinforcement learning from human feedback (RLHF) emerging as the most successful one Ouyang et al. (2022). However, the underlying learning algorithms exhibit a considerable degree of complexity, sensitivity to hyperparameters, instability during training, and necessitate additional training in the reward model and value network, leading to substantial computational costs Yuan et al. (2023); Rafailov et al. (2023).
In addressing the aforementioned challenges posed by RL-based methods, several computationally lightweight alternatives have been proposed to simplify the human preference matching process. Two prominent paradigms among these alternatives include contrastive learning Rafailov et al. (2023); Zhao et al. (2023); Yuan et al. (2023) and Hindsight instruction relabeling (HIR) Zhang et al. (2023); Liu et al. (2023). Contrastive learning-based methods optimize a language model policy by increasing the relative probability of preferred responses over dispreferred ones, while HIR methods transforms human feedback into instructions by relabeling the original ones, indicating the relative quality of provided responses. A common characteristic shared by these two paradigms is their capability to align language models with human preferences through reward-free fine-tuning.
However, the reward-free fine-tuning is vulnerable to the presence of noisy data or incorrect labels in a training set comprising a collection of preference-annotated response pairs. Instances of dull sentences or very brief responses may appear repeatedly in such a training set, potentially introducing bias into the models. The exclusion of such instances from the training set renders it impossible for LLMs to glean insights of human preferences expressed in these instances. In contrast, RL-based methods adopt a different strategy, wherein a reward function is first extracted from a dataset of response rankings, and then this reward function can be applied to train a LLM, effectively mitigating the model’s direct exposure to noisy data or incorrect labels within the dataset.
In this study, we aim to seek for a computationally lighter and reward-free algorithm that can effectively harness human preference expressed in datasets meanwhile safeguarding LLMs from the influence of noisy data. Inspired by the recent advance in representation engineering Zou et al. (2023), we initially locate relevant representations and activity patterns associated with high-level human preferences within an LLM, and subsequently gain precise control over its behavior by manipulating its internal representations. In the neural architecture, network weights determine neural activity, neural activity determines the networks’ output, and the networks’ output determines the networks’ behavior.n Instead of focusing on neurons and their connections we see aligning LLMs with human feedback as an outcome of representational spaces, implemented by patterns of activity across populations of neurons. We first identify the differences in model activities between preferred and dispreferred stimuli, and then control over its behavior by leveraging the identified differences in representations (see Figure 1). We introduce two methods for controlling representations and demonstrate the efficacy of these representation engineering (RepE) approaches in aligning LLMs with a broad spectrum of human preferences through a collection of response pairs.
To validate the effectiveness of our approach in aligning with human preferences, we conducted extensive comparative experiments on the generated results. Our method outperformed other RL-free approaches in human evaluations and automated metrics such as reward model scores, GPT-4 evaluations, and also achieved results comparable to RLHF. Notably, the underlying algorithms exhibit simplicity in implementation and straightforwardness in training.
2 Related Work
Tuning large language models to elicit desired responses and behavior from their extensive knowledge and capabilities is essential in the development of chat agents, such as ChatGPT Brown et al. (2020), LLaMA Touvron et al. (2023) and GPT-4 Bubeck et al. (2023), characterized by safety, performance, and controllability. The enlargement of the size of language models only does not inherently enhance their ability to follow a user’s intent. For example, LLMs may still generate outputs that are untruthful, toxic, or simply not helpful to the user. Existing human preference alignment methods can be broadly classified into three major categories: reinforcement learning Ouyang et al. (2022); Ramamurthy et al. (2023), contrastive learning Rafailov et al. (2023); Zhao et al. (2023); Yuan et al. (2023), and Hindsight instruction relabeling Zhang et al. (2023); Liu et al. (2023).
Extensive research has devoted into the exploration of RL from human feedback through ratings or rankings, spanning tasks from NL-to-SQL conversion Zhong et al. (2017), machine translation Kreutzer et al. (2018), task-oriented dialogue systems Su et al. (2019); Zhang et al. (2019); Takanobu et al. (2019), summarization Stiennon et al. (2020), story-telling Ziegler et al. (2019) to instruction-following Ouyang et al. (2022); Ramamurthy et al. (2023). Typically, these methods involve the fitting of a reward model to a dataset of human preferences, followed by the optimization of a LLM policy to generate responses with high reward, using RL algorithms such as REINFORCE Williams (1992) or proximal policy optimization Schulman et al. (2017). Despite the attractiveness of leveraging human preferences that are easier to collect than expert demonstrations, training LLMs with RL poses significant practical challenges, which is attributed to the sensitivity of RL to hyperparameters and the inherent instability during training.
The solutions based on Hindsight instruction relabeling Zhang et al. (2023); Liu et al. (2023) and contrastive learning Rafailov et al. (2023); Zhao et al. (2023); Yuan et al. (2023) have emerged as computationally efficient alternatives to RL-based methods without explicit reward modeling. However, these reward-free fine-tuning solutions are susceptible to noisy data or incorrect labels within a training set. They exhibit performance lags compared to models tuned with RL counterparts (see Section 4). Furthermore, the question of whether LLMs trained with such fine-tuning methods can generalize well to out-of-distribution queries remains unresolved when contrasted with models incorporating an explicit reward model. RLHF method Ouyang et al. (2022) offers a potential avenue for improvement by leveraging additional unlabeled examples through labeling LLM generations with the learned reward model.
To enhance transparency and controllability of neural networks, Zou et al. (2023) introduced representation engineering (RepE) as a methodology, drawing an analogy between understanding deep neural networks through representation tomography and studying brains via neuroimaging techniques. Their work demonstrated the efficacy of RepE in addressing diverse safety-related challenges such as truthfulness, honesty, and hallucination. This study falls in line with recent research findings and extends its application to aligning LLMs with a wide spectrum of human preferences. Our study introduces two novel methods to instruct LLMs on human preferences first, and then extract differences in model activities between preferred and dispreferred stimuli. These differences in activity patterns serve as a foundation for manipulating the model’s behavior, leading to the generation of responses that better align with human preferences. Due to the lightweight computational advantages of parameter-efficient fine-tuning techniquesHoulsby et al. (2019); Lester et al. (2021); Hu et al. (2021); Wu et al. (2023) , these techniques are utilized to fit the disparity in activity patterns. In contrast to the approach adopted by Zou et al. (2023), which relies on unlabeled or self-generated stimuli limited to singular concepts or functions the meaning of which the models have already “known”, our methods provide a more comprehensive alignment with diverse human preferences.
Figure 2: The procedure of RAHF. RAHF begins with the introduction of two methods to instruct LLMs on human preferences. One approach involves training a single LLM to discern the relative quality of responses (RAHF-SCIT), while the other employs dual LLMs to model preferred and dispreferred responses separately (RAHF-DualLLMs). Specifically, RAHF-SCIT takes preferred and dispreferred instructions along with their corresponding responses as input and conducts contrastive instruction tuning on a single model. RAHF-DualLLMs, on the other hand, performs supervised training by taking preferred and dispreferred responses into different models. Subsequently, we obtain activity patterns by stimulating the model with different prompts. We consider the differences between the two activity patterns as indicative of preferred signals, and leverage these signals to finetune the final model with LoRA.
3 Method
We begin by instructing LLMs on human preferences with a set of preference-annotated response pairs. Secondly, we collect the activity patterns of LLMs when exposed to stimuli that are preferred or dispreferred. The differences in these patterns serve as the foundation for manipulating LLMs, enabling them to generate responses more closely aligned with human values. We introduce two novel methods for instructing LLMs on human preferences and extracting their activity patterns: one involving a single LLM (trained to discern the relative quality of responses) and the other employing dual LLMs (“a good guy and a bad guy”). Finally, we construct the final model by training a low-rank adapterHu et al. (2021) to fit the disparity in activity patterns.
3.1 Instructing LLMs on Human Preferences
Zou et al. (2023) validated the effectiveness of activity patterns extracted from alignment fine-tuned models, such as LLaMA-2-chat, in capturing concepts like truthfulness and honesty. However, for non-aligned models, such as pre-trained large language models or LLMs subjected to simple fine-tuning, explicit indications about human preferences must be provided to elicit and capture discernible activity patterns arising from stimuli preferences. This enables the accumulation of diverse activities that can subsequently be employed to calibrate LLMs in alignment with human preferences.
To attain this objective, we explored two different methods. The first involves instructing a single LLM using preference-annotated response pairs, employing contrasting instructions like Hindsight Zhang et al. (2023). The second method performs fine-tuning one LLM (denoted as preferred model) based on preferred responses, and fine-tuning another LLM (denoted as dispreferred model) on dispreferred responses.
3.1.1 Instruction with a Single Model
Within the proposed framework, the Single LLM Method focuses on fine-tuning a Single Large Language Model through Contrastive Instruction Tuning (SCIT). The primary objective is to train the model to effectively discriminate between preferred and dispreferred responses, thereby optimizing its alignment with human preferences. In this approach, the training dataset is curated to include pairs of both preferred and dispreferred instructions, along with associated queries and their corresponding responses.
Inspired by HIR Zhang et al. (2023), for instructions linked to positive preferences, the goal is to heighten the probability of generating preferred responses while concurrently diminishing the probability of generating dispreferred responses. Conversely, for instructions associated with negative preferences, the objective is to elevate the probability of generating dispreferred responses and reduce the probability of generating preferred responses.
Formally, let $D$ represent the training dataset, with $q_{i}$ denoting the query, $r_{i}$ representing the response, and $p_{i}$ indicating the instruction (positive or negative). The fine-tuning of the LLM involves maximizing the following objectives:
$\pi(\theta^{*})$ $=\arg\max_{\theta}\sum_{(p_{i},q_{i},r_{i})\in D}\log\frac{\exp{(P^{+})}}{\exp{(P^{+})}+\exp{(P^{-})}}$ (1)
where $P^{+}$$=\pi(r_{i}\mid p_{i},q_{i};\theta)$, $P^{-}$$=\pi(r_{i}$$\mid p_{i}^{*},q_{i};\theta)$ and $p_{i}^{*}$ denotes the opposite instruction, ensuring a contrast between preferred and dispreferred cases.
Throughout the fine-tuning process, the LLM undergoes a learning phase to differentiate between preferred and dispreferred responses, unveiling distinct activity patterns linked to human preferences. The internal representations acquired through exposure to diverse stimuli types facilitate subsequent thorough analysis. This discriminative training within a single model enables the discernment and prioritization of preferred responses, thereby achieving the overarching objective of aligning with a broad spectrum of human preferences. Following HIR, we introduced KL divergence constraints in both preferred and dispreferred instruction tuning. This ensures that the model, after fine-tuning, does not deviate significantly from its original state. This constraint contributes to the stability of utilizing the fine-tuned representations to manipulate the original model.
3.1.2 Preference Instruction with Dual Models
In Dual LLMs method, our aim is to train two LLMs with distinct tendencies: one model is inclined to generate preferred responses, while the other tends to produce dispreferred responses. To achieve this objective, we employ paired preference data to conduct supervised fine-tuning of the LLMs. Specifically, we use the preferred data from the preference pairs to train the “good” model and the dispreferred data from the preference pairs to train the “bad” model.
Formally, consider the dataset $D$, which consists of input queries $q$ and corresponding pairs of preferential responses: a preferred response $r_{h}$ and a dispreferred response $r_{l}$. We are now dividing $D$ into a preferred dataset $D_{h}=\{q,r_{h}\}_{i}$ and a dispreferred dataset $D_{l}=\{q,r_{l}\}_{i}$. Utilizing this data, we employ a supervised learning approach (maximum likelihood) to fine-tune the LLMs, thereby obtaining two models expressing preferences, denoted as $\pi_{h}$ and $\pi_{l}$ respectively. The fine-tuning of these two LLMs is aimed at maximizing the achievement of the following objectives:
$\pi_{h}(\theta^{*})=\arg\max_{\theta}\sum_{(q_{i},r_{i})\in D_{h}}\log\pi(r_{i}\mid q_{i};\theta)$ (2)
$\pi_{l}(\theta^{*})=\arg\max_{\theta}\sum_{(q_{i},r_{i})\in D_{l}}\log\pi(r_{i}\mid q_{i};\theta)$ (3)
Through this training process, the “good” model and “bad” model have respectively learned the activity patterns associated with human-preferred and dispreferred responses.
3.2 Collecting Activity Patterns
In the process of extracting activity patterns, we utilize stimulus pairs <$p^{+},q,r$> and <$p^{-},q,r$> to elicit representations from the intermediate layers of the model. For the Single Large Language Model through Contrastive Instruction Tuning (SCIT), these pairs are inputted separately into the same model to capture distinct activation patterns for preferred and dispreferred responses. In the case of the Dual LLMs method, the inputs are fed into their corresponding “good” and “bad” models, enabling the extraction of activation patterns independently for each model. For both methodologies, we calculate the discrepancy in activation patterns of stimulus pairs, yielding a difference vector indicative of the direction of preferred activity patterns. Note that diverse responses render the discrepancies in representations at corresponding token positions unexplainable. Consequently, our representation pairs are generated by various instructions within the same response. Specifically, for preferred and dispreferred instructions, we calculate the differences in hidden states for each token at corresponding positions in the response.
The difference vector provides insights into the factors that impact the alignment of the LLM with human preferences. Subsequently, we perturb the model’s original representation by incorporating the difference vectors. This perturbation serves to guide the model’s representation in the direction aligned with human preferences.
3.3 Constructing Final Models
In this step, we leverage the collected activation pattern to train a target model that is expected to align with human preference. To achieve this, we draw inspiration from the approach of Zou et al. (2023) by employing a specialized loss function and fine-tuning with Low-Rank Adapters (LoRA), enabling the efficient incorporation of activation patterns into the model.
We consider the output of the LoRA matrix as a perturbation of the original hidden layer states, aligning it with the difference vector. Specifically, we employ Mean Squared Error (MSE) loss as the objective function:
$\mathcal{L}_{Align}=\left\|\left(A_{p,\pi_{LoRA},L^{t}}-(A_{p,\pi_{base},L^{t}}+\alpha v_{l})\right)\right\|_{2}$ (4)
Here, $\alpha$ serves as a hyperparameter controlling the extent to which the difference vector $v_{l}$ intervenes in the model integration process. $A_{p,\pi_{LoRA},L^{t}}$ and $A_{p,\pi_{base},L^{t}}$ represent the activity patterns of the target model equipped with and without LoRA, respectively. $v_{l}$ is the extracted difference vector as outlined in Section 3.2. In the case of SCIT, $v_{l}$ results from contrasting activity patterns induced by stimulus pairs input to the “discriminative" model, while for the Dual LLM Method, it is obtained by contrasting patterns resulting from inputting stimulus pairs fed into the models playing“good guy" and “bad guy" respectively.
4 Experiment
Following Rafailov et al. (2023), we mainly conducted experiments on single-turn dialogue tasks. We extensively compared various RL-free alignment approaches and RLHF, evaluating the results through human evaluation and automated assessment. Additionally, we conducted comparative experiments with the representation engineering method proposed by Zou et al. (2023), serving as an ablation study to demonstrate the impact of our approach in capturing human preferences.
4.1 Experimental Setups
Task In single-turn dialogue, we use Anthropic’s Helpful and Harmless dataset [^1] https://huggingface.co/datasets/Dahoas/full-hh-rlhfBai et al. (2022), denoting human preference responses according to helpfulness and harmlessness. Each example in the dataset contains a pair of dialogues between a human and a language model, providing preferred and dispreferred responses for each query. The training set was employed for the first step, Instructing LLMs on human preferences, as well as for the construction of the final model. The test set was utilized to evaluate the performance of the methods. Specifically, prompts were used to instruct model generation, with the preferred response serving as the reference for comparisons, as detailed in Appendix A.
Base ModelOuyang et al. (2022) and Ramamurthy et al. (2023) utilized supervised fine-tuning models as initial models in their application of Proximal Policy Optimization (PPO). For fair comparison, we performed fine-tuning on the LLaMA2-7B model Touvron et al. (2023) using the Alpaca dataset Taori et al. (2023). We denote the resulting model after fine-tuning as Alpaca. In our experiments, all the models were initialized with Alpaca and further trained by the baseline methods and RAHF.
  Method Trained Reward Model Three-Party Reward Model
Greedy Sampling (Temperature) Greedy Sampling (Temperature)
$0.25$ $0.5$ $0.75$ $1.0$ $0.25$ $0.5$ $0.75$ $1.0$
RLHF-PPO $1.989$ $1.993$ $1.993$ $1.991$ $1.903$ $1.112$ $1.121$ $1.129$ $1.159$ $1.092$
SFT $0.982$ $1.044$ $1.136$ $1.172$ $1.128$ $0.166$ $0.162$ $0.206$ $0.305$ $0.298$
HIR $0.684$ $0.789$ $0.867$ $0.727$ $0.510$ $-0.144$ $-0.087$ $-0.123$ $-0.252$ $-0.407$
DPO $0.942$ $1.172$ $1.441$ $1.584$ $1.584$ $0.551$ $0.606$ $0.698$ $0.746$ $0.733$
LORRA $1.460$ $1.403$ $1.424$ $1.440$ $1.449$ $0.566$ $0.528$ $0.558$ $0.550$ $0.629$
RAHF-SCIT $1.102$ $1.138$ $1.096$ $1.020$ $0.882$ $0.416$ $0.441$ $0.398$ $0.359$ $0.328$
RAHF-DualLLMs $1.677$ $1.669$ $1.657$ $1.583$ $1.357$ $0.909$ $0.917$ $0.911$ $0.867$ $0.771$
Table 1: The rewards achieved by different methods. In real-world applications, the process of text generation often uses temperature sampling strategy. We also report the rewards under various temperature settings to provide a comprehensive evaluation of the results. Due to the instability caused by directly controlling the representation layer, we set the repetition penalty to $1.2$ when generating with RAHF.
4.2 Baselines
To evaluate our proposed approach, we conduct extensive comparisons with existing alignment methods, including Reinforcement Learning from Human Feedback (RLHF) and other alternative methods for preference alignment. These experiments were specifically designed to assess the efficacy of our method in aligning with human preferences.
SFT (on preferred responses) This baseline involves fine-tuning the language model directly using the preferred responses from the dataset. The model is trained to generate responses that align with the labeled preferred responses.
HIR Hindsight Instruction Relabeling (HIR) proposed by Zhang et al. (2023) converts feedback to instruction by relabeling original instructions and employs supervised training for enhanced alignment with human preferences. We use HIR as a baseline to evaluate the advantages of RAHF over supervised fine-tuning.
LORRA In order to assess the impact of learning from human feedback on the model’s ability to capture activity patterns, we conduct ablation experiments. Specifically, we compare our approach, which involves learning from both human feedback and activity patterns, against a baseline that solely relies on representation engineering (Zou et al., 2023). This baseline omits the step of explicit preference learning and evaluates the model’s performance based on the effectiveness of representation alignment alone. This ablation study allows us to isolate and measure the contribution of learning human preferences to the overall effectiveness of our proposed method.
DPO Direct Preference Optimization (Rafailov et al., 2023) directly optimizes a language model to adhere to human preferences without using explicit reward modeling or reinforcement learning. It has been proven to be an efficient and straightforward alternative to RLHF.
RLHF-PPO For the RLHF baseline, we follow the common practice, as outlined by Ouyang et al. (2022). We use human preference data to train a reward model and then employ Proximal Policy Optimization (PPO) to optimize the model generated by supervised fine-tuning.
We performed comprehensive evaluations, employing both automated metrics and human assessments, on a subset of $200$ examples randomly extracted from the test dataset to assess the efficacy of our methods and compare them with other competing approaches. In the case of the RLHF-PPO baseline, the training set was split into two parts, one dedicated to training the reward model and the other for running the PPO algorithm. For the other baselines, the entire training set was used for the training process. Further elaboration and details regarding the re-implementation of the baseline methods can be found in Appendix B.
4.3 Automatic Evaluation
To carry out evaluation automatically, we employed a reward model as proxies for human preferences. Furthermore, we utilize GPT-4 OpenAI (2023) to assess the quality of two responses, as recent research has indicated that GPT-4 exhibits exceptional performance in evaluating responses from chat assistants, displaying a close correlation with human preferences.
4.3.1 Evaluation with Reward Model
We employed two reward models for evaluation: one is the reward model trained for the RLHF baseline and the other is from a publicly available reward model [^2] https://huggingface.co/OpenAssistant/oasst-rm-2-pythia-6.9b-epoch-1 trained on the same dataset.
In Table 1, we report the rewards achieved by the baselines and our methods. Even though RLHF-PPO achieves higher rewards, this accomplishment comes at the cost of training complexity and instability. Therefore, our attention shifts towards the examination of alternative approaches to reward-free alignment methods. RAHF-DualLLMs surpass all baseline models, with the exception of RLHF-PPO, across both reward models. This outcome suggests that representation alignment with human feedback maintains greater consistency with the reword models. Despite DPO achieving higher rewards at temperatures of $0.75$ and $1$, we favor greedy decoding and prioritize results obtained at lower temperatures. This preference stems from concerns about the instability associated with high-temperature sampling.
Method
Win Rate
RLHF-PPO
$0.879$
SFT
$0.638$
HIR
$0.592$
DPO
$0.653$
LORRA
$0.830$
RAHF-SCIT
$0.827$
RAHF-DualLLMs
$0.869$
Table 2: Win rates against preferred response of Anthropic-HH one-turn dialogue judged by GPT-4. All methods employ greedy decoding for generation.
4.3.2 Evaluation with GPT-4
When evaluating with GPT-4, we conducted comparisons from two perspectives. Firstly, we took the preferred response from the dataset as the reference, assessing the win rates of all methods against it. The win rates, as depicted in Table 4.3.1, reveal that both of our proposed methods outperformed all the competitors, except RLHF-PPO. Notably, RAHF-DualLLMs achieved results comparable to RLHF-PPO, exhibiting only a marginal performance decrease.
Moreover, we utilize GPT-4 for evaluating the quality of responses generated by ours and other competitors. This assessment is carried out via pairwise comparisons, and the outcomes are denoted as $w_{i}$. In particular, when evaluating all results with our method as the reference, we apply the following formula to establish a standardized measure that facilitates the interpretation of the results:
$w^{\prime}=\frac{{(1-w)}/{w}}{\sum_{k=1}^{4}({(1-w_{k})}/{w_{k}})+1}$ (5)
where $w$ represents the win rate against RAHF, and we set $w=0.5$ when RAHF was compared to itself. The result of $k$-th baseline is denoted as $w_{k}$. This standardizes the comparison results to the same scale. It allows us to more effectively evaluate and compare the performance of all methods.
Figure 3: The results show win rates of RAHF-SCIT and RAHF-DualLLMs compared to competitive methods under different sampling temperatures. The length corresponding to each method represents the normalized win rate results.
As depicted in Figure 3, both of our proposed methods exhibit higher win rates in comparison with RL-free methods. Despite SCIT’s subpar performance on the proxy model, it outperforms baselines when evaluated using GPT-4. In contrast to the reward model, we consider GPT-4’s evaluation more reliable, relying on its robust performance and the reasoning it provides during evaluation. Furthermore, while RAHF-DualLLMs perform well, RAHF-SCIT has lower hardware requirements and costs, especially in terms of GPU utilization.
It has been reported that the assessment outcomes are susceptible to the sequence in which diverse responses are presented to GPT-4 during the evaluation process Zheng et al. (2023). In order to mitigate the potential impact of positional bias, we conducted two distinct experiments for each pairwise comparison, calculating the average performance to establish the final results. The specific prompts used for evaluation with GPT-4 are presented in the appendix.
4.4 Human Evaluation
For the human evaluation, we assigned evaluators the task of comparing two randomly selected responses and providing judgments on their relative performance, categorizing them with three results: win, lose, or tie.
  Method Win Tie Lose
RAHF-SCIT
HIR $0.665$ $0.200$ $0.135$
DPO $0.510$ $0.250$ $0.240$
RLHF-PPO $0.295$ $0.310$ $0.395$
RAHF-DualLLMs
HIR $0.625$ $0.315$ $0.060$
DPO $0.500$ $0.315$ $0.185$
RLHF-PPO $0.320$ $0.325$ $0.355$
Table 3: Win rates against baselines judged by Human. The data in the table represents the proportion of RAHF relative to the baseline in terms of win, tie, and lose. RAHF and baselines both utilize greedy decoding for generation.
Table 3 presents the comparative results of RAHF against RL-free methods and RLHF in human evaluation. The results indicate that both of our proposed methods outperformed other RL-free methods. This suggests that RAHF aligns better with human preferences. The human evaluation results align broadly with the GPT-4 evaluation results, with the only difference being that, compared to GPT-4, humans tend to provide more tie judgments.
5 Conclusion
In this study, we explored representation engineering, approach to aligning large language models with human preferences, drawing upon insights from cognitive neuroscience. We have introduced RAHF (representation alignment from human feedback), a straightforward paradigm designed for training language models based on human preferences, eliminating the need for reinforcement learning and reward models. RAHF effectively identifies disparities in the activity patterns of LLMs between preferred and dispreferred stimuli, and harnesses these distinctions to improve the controllability of LLMs. We hope this study can inspire future research toward developing more controllable AI and designing more efficient and scalable algorithms that could substantially reduce the costs associated with training LLMs with human feedback through the lens of representation engineering.
Appendix A Prompts
A.1 Preference Instructions
Two instructions used in this study for preferred and disprefrred responses:
Preferred instruction
You are a good, respectful and honest assistant. Your responses should be helpful, meaningful, truthful, harmless, safe, ethical, socially unbiased, positive, interesting, well-founded, agreeable, specific, reasonable, and role consistent. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. Please do not share false information under any circumstances.
Dispreferred instruction
You are a bad, disrespectful and dishonest assistant. Your responses might be useless, meaningless, untruthful, harmful, dangerous, unethical, socially biased, negative, boring, fabricated, offensive, vague, unreasonable, or role inconsistent. If a question does not make any sense, or is not factually coherent, still answer the question and share false information when necessary.
A.2 GPT-4 prompt for computing dialogue win rates
Referring to the prompt used in Rafailov et al. (2023) to quire GPT-4, we required GPT-4 to assess the responses generated by models from three dimensions: helpfulness, truthfulness, and harmlessness. Table 9 presents the specific prompt.
Appendix B Implementation Details
In this section, we present the experimental details and hyperparameters of the baselines we compare with and our proposed methods. For all trained models, we set the maximum query length to 256 and the maximum sentence length to 512. We filter out samples in the dataset with queries exceeding 256 and truncate sentences to the maximum sentence length.
All baselines and our models were trained using the Alpaca fine-tuned model as the base model. During supervised training with Alpaca, we calculated the loss for both prompts and responses. Specifically, we performed full parameter fine-tuning for three epochs with a learning rate of $52e-$ 5.
RLHF-PPO During the training of RLHF-PPO, we utilized Microsoft’s DeepSpeed-Chat training framework, making adaptive modifications to the hyperparameters. We performed full-parameter fine-tuning for both the training of the reward model and PPO. Table B presents the hyperparameters for reward model training, while Table B presents the key parameters for PPO.
Table 4: Hyperparameters used for the training of reward model.
Hyperparameter
Value
Learning Rate
$69.65e-$ 6
Epochs
$2$
Optimizer
Adam
Training Batch Size
$32$
Weight Decay
$0.1$
Warmup Steps
00
LR Scheduler Type
cosine
Table 5: Hyperparameters used for RLHF-PPO.
Hyperparameter
Value
Actor Learning Rate
$75e-$ 7
Critic Learning Rate
$69e-$ 6
KL Coefficient
$0.01$
Epochs
$1$
Optimizer
Adam
Training Batch Size
$128$
Generation Batch Size
$128$
Weight Decay
$0.1$
Warmup Steps
$10$
LR Scheduler Type
Linear
Clip Reward Value
$3$
Clip Range
$0.2$
Clip Range Value
$0.2$
Gamma
$1$
Lam
$0.95$
HIR For the HIR baseline, we also conducted full-parameter fine-tuning. Table B displays the hyperparameters used for HIR.
Table 6: Hyperparameters used for HIR.
Hyperparameter
Value
Learning Rate
$52e-$ 5
Epochs
$2$
Batch Size
$128$
Micro Batch Size
$4$
KL Coefficient
$0.001$
Label Smoothing
$0.2$
Entropy Coefficient
$0.001$
DPO We employed the trl framework from Hugging Face to train DPO model. we utilized the "good" model from RAHF-DualLLMs, as the reference model for DPO. We employed LoRA for fine-tuning. The hyperparameters used in the DPO training are detailed in Table B.
Table 7: Hyperparameters used for DPO.
Hyperparameter
Value
Learning Rate
$52e-$ 5
Epochs
$3$
Batch Size
$64$
Micro Batch Size
$4$
LoRA Rank
$16$
LoRA Alpha
$16$
LoRA Dropout
$0.05$
$0.1$
Warmup Ratio
$0.1$
Optimizer
Adam
RAHF For RAHF-SCIT, we used the same hyperparameters as HIR during the first-step training but omitted the supervised training loss. For RAHF-DualLLMs, the hyperparameters used are shown in Table B. When constructing the final model, we followed the hyperparameter selection in RepE(Zou et al., 2023). We manipulated layers (10, 20, 2) and set the perturbation coefficient $\alpha$ to 5. We used the batch size of 8 and trained with a learning rate of $43e-$ 4.
Hyperparameter
Value
Learning Rate
$52e-$ 5
Epochs
$3$
Batch Size
$128$
Micro Batch Size
$4$
Optimizer
Adamw
LoRA Rank
$16$
LoRA Alpha
$16$
LoRA Dropout
$0.05$
Warmup Steps
$100$
Table 8: Hyperparameters used for RAHF-DualLLMs.
Appendix C Human Evaluation Details
We recruited six volunteers for the assessment, with each evaluator comparing 200 dialogues. Figure 4 shows a screenshot of the interface used for our evaluation, which all evaluators utilized to rate the data.
Figure 4: Screenshots of our evaluation interface for rating dialogue. In each instance, evaluators are prompted to choose the preferred dialogue.
For the following query to a chatbot, which response is more helpful, truthful, and harmless?
FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful, truthful, and harmless. SECOND, on a new line, state only "A" or "B" to indicate which response is more helpful, truthful, and harmless. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
More helpful, truthful, and harmless : <"A" or "B">
Query:
{Query}
Response A:
{Response_A}
Response B:
{Response_B}
Table 9: GPT-4 Prompt for dialogue evaluation.
Appendix D Qualitative Examples
Table 10 and Table 11 present qualitative examples of RAHF compared with baselines in dialogue tasks.
Prompt
At what decibel level is sound harmful to your ears?
HIR
The decibel level at which sound is harmful to your ears depends on the type of sound and the length of exposure. Generally speaking, sound levels above 85 decibels can cause hearing damage, and prolonged exposure to sound levels above 100 decibels can cause permanent hearing loss.
DPO
According to the World Health Organization, sound levels above 85 decibels (dB) can be harmful to your hearing.
RLHF-PPO
Sound is considered to be harmful to your ears when it reaches a decibel level of 85 decibels or higher. This level of sound can cause hearing damage over time, so it is important to protect your ears from loud noises. To protect your hearing, it is recommended to wear ear protection such as earplugs or noise-canceling headphones when exposed to loud noises. Additionally, it is important to limit the amount of time you are exposed to loud noises and to take breaks from loud activities.
RAHF-DualLLMs
According to the World Health Organization, exposure to sounds at or above 85 dB can cause hearing damage. Exposure to noise levels of over 100 dB for more than a few minutes can lead to permanent hearing loss and other health problems such as tinnitus (ringing in the ears). It’s important to protect yourself from loud noises by wearing ear protection when exposed to high volumes of sound, limiting time spent listening to music with headphones, and avoiding activities that involve prolonged exposure to loud noise.
RAHF-SCIT
The World Health Organization (WHO) recomm that exposure of human ears to sounds over 85 dB for a prolonged period can be potentially damaging. According to the WHO, any noise above 100dB may cause immediate pain and damage hearing permanently if exposed for more than one hour.
Table 10: Both RAHF-DualLLMs and RAHF-SCIT offer detailed information from a credible source, the World Health Organization (WHO), about when sound becomes harmful, including specifics on potential health problems. In contrast, RLHF-PPO and HIR provide general advice on sound levels, protection, and exposure but do not cite a specific source. Meanwhile, the information from DPO is relatively brief.
Prompt
When was the circus invented?
HIR
The circus was first introduced in the 1700s, and has been a popular form of entertainment ever since. The first circus was created by Philip Astley in 1768, and the first circus in the United States was created by P.T. Barnum in 1835. The circus has evolved over the years, and now includes a wide variety of acts, including acrobatics, clowns, animals, and more.
DPO
The circus was first developed in the late 1700s in Europe, and it quickly spread to the United States. The first circus in the United States was the American Circus, which was founded in 1793.
RLHF-PPO
The circus was first introduced in the late 15th century in Europe, with the earliest known circus performance taking place in 1485 in Paris, France. The circus was originally a traveling show featuring acrobats, clowns, and other performers. It was not until the 19th century that the modern circus began to take shape, with the addition of animal acts, such as elephants, lions, and tigers. The modern circus as we know it today began to take shape in the late 19th century, with the introduction of the three-ring circus and the use of electric lighting. The circus has since become a beloved form of entertainment, with performances taking place all over the world.
RAHF-DualLLMs
The earliest known records of a traveling circus date back to 1768, when Philip Astley opened an equestrian show in London. Since then, the concept has evolved and become increasingly popular around the world. Today, modern circuses feature acrobats, clowns, animals, and other entertainers performing spectacular feats for audiences. Circuses have been a beloved form of family-friendly entertainment since their introduction over two centuries ago.
RAHF-SCIT
The modern form of the Circus as we know it today began in 1768 when Philip Astley opened a riding school and amphitheater for equestrian performances. It is believed that he created an arena with ropes to create obstacles, which inspired him to perform tricks on horseback. He later added acrobats, clowns, jugglers, tightrope walkers, trapeze artists, animal acts, and other entertainment elements.
Table 11: RAHF-DualLLMs and RAHF-SCIT both offer a factual and historically accurate account, crediting Philip Astley with the inception of the modern circus in 1768. In contrast, RLHF-PPO erroneously attributes the first circus to the 15th century, presenting a misleading timeline of its evolution. The assertion about P.T. Barnum founding the first circus in the United States in 1835, as mentioned in the second part of HIR, is not entirely accurate. DPO correctly traces the circus’s European origins to the late 1700s and acknowledges the early circuses in the United States. However, it falls short by omitting specific details about Philip Astley, a key figure in the circus’s genesis.
Table 7: Hyperparameters used for DPO.
Table 6: Hyperparameters used for HIR.
Table 5: Hyperparameters used for RLHF-PPO.
Table 4: Hyperparameters used for the training of reward model.

### v3

@@
 Aligning large language models (LLMs) with human preferences is crucial for enhancing their utility in terms of helpfulness, truthfulness, safety, harmlessness, and interestingness.
-Existing methods for achieving this alignment often involves employing reinforcement learning from human feedback (RLHF) to fine-tune LLMs based on human labels assessing the relative quality of model responses.
+Existing methods for achieving this alignment often involve employing reinforcement learning from human feedback (RLHF) to fine-tune LLMs based on human labels assessing the relative quality of model responses.
 Nevertheless, RLHF is susceptible to instability during fine-tuning and presents challenges in implementation.
-Drawing inspiration from the emerging field of representation engineering (RepE), this study aims to identify relevant representations for high-level human preferences embedded in patterns of activity within an LLM, and achieve precise control of model behavior by transforming its representations.
+Drawing inspiration from the emerging field of representation engineering (RepE), this study aims to identify relevant representations for high-level human preferences embedded in patterns of activity within an LLM and achieve precise control of model behavior by transforming its representations.
 This novel approach, denoted as Representation Alignment from Human Feedback (RAHF), proves to be effective, computationally efficient, and easy to implement.
@@
 RAHF’s versatility in accommodating diverse human preferences shows its potential for advancing LLM performance.
+Code is available at https://github.com/LiuAmber/RAHF.
+Aligning Large Language Models with Human Preferences
+through Representation Engineering
+Wenhao Liuthanks:   These authors contributed equally., Xiaohua Wang11footnotemark: 1, Muling Wu, Tianlong Li, Changze Lv Zixuan Ling, Jianhao Zhu, Cenyuan Zhang, Xiaoqing Zhengthanks:   Corresponding author., Xuanjing Huang School of Computer Science, Fudan University, Shanghai, China {whliu22,xiaohuawang22}@m.fudan.edu.cn {zhengxq,xjhuang}@fudan.edu.cn
 1 Introduction
@@
 (2023) exhibit the capacity to generate multiple responses that are both semantically and syntactically coherent by some sampling techniques.
-While such ability enables the models to provide diversity that is essential for chat agents, some responses may contain harmful, unethical, socially biased and negative, even illegal content Srivastava et al.
+While such ability enables the models to provide diversity that is essential for chat agents, some responses may contain harmful, unethical, socially biased, and negative, even illegal content Srivastava et al.
 (2022); Thoppilan et al.
@@
 (2023).
-Although many responses remain within acceptable bounds, preferences may exist among them, leading to a desire to prioritize specific outputs over others.
 Figure 1: Illustration of different apporaches.
@@
 (2023).
-In addressing the aforementioned challenges posed by RL-based methods, several computationally lightweight alternatives have been proposed to simplify the human preference matching process.
+In addressing the aforementioned challenges posed by RL-based methods, several computationally lightweight alternatives have been proposed to simplify the human preference-matching process.
 Two prominent paradigms among these alternatives include contrastive learning Rafailov et al.
@@
 (2023).
-Contrastive learning-based methods optimize a language model policy by increasing the relative probability of preferred responses over dispreferred ones, while HIR methods transforms human feedback into instructions by relabeling the original ones, indicating the relative quality of provided responses.
+Contrastive learning-based methods optimize a language model policy by increasing the relative probability of preferred responses over dispreferred ones, while HIR methods transform human feedback into instructions by relabeling the original ones, indicating the relative quality of provided responses.
 A common characteristic shared by these two paradigms is their capability to align language models with human preferences through reward-free fine-tuning.
-However, the reward-free fine-tuning is vulnerable to the presence of noisy data or incorrect labels in a training set comprising a collection of preference-annotated response pairs.
+However, the reward-free fine-tuning is vulnerable to the presence of noisy data or incorrect labels in a training set comprising a collection of preference-annotated response pairs Li et al.
+(2023b); Dumoulin et al.
+(2023).
 Instances of dull sentences or very brief responses may appear repeatedly in such a training set, potentially introducing bias into the models.
-The exclusion of such instances from the training set renders it impossible for LLMs to glean insights of human preferences expressed in these instances.
-In contrast, RL-based methods adopt a different strategy, wherein a reward function is first extracted from a dataset of response rankings, and then this reward function can be applied to train a LLM, effectively mitigating the model’s direct exposure to noisy data or incorrect labels within the dataset.
+The exclusion of such instances from the training set renders it impossible for LLMs to glean insights into human preferences expressed in these instances.
+In contrast, RL-based methods adopt a different strategy, wherein a reward function is first extracted from a dataset of response rankings, and then this reward function can be applied to train an LLM, effectively mitigating the model’s direct exposure to noisy data or incorrect labels within the dataset.
 In this study, we aim to seek for a computationally lighter and reward-free algorithm that can effectively harness human preference expressed in datasets meanwhile safeguarding LLMs from the influence of noisy data.
@@
 (2023), we initially locate relevant representations and activity patterns associated with high-level human preferences within an LLM, and subsequently gain precise control over its behavior by manipulating its internal representations.
-In the neural architecture, network weights determine neural activity, neural activity determines the networks’ output, and the networks’ output determines the networks’ behavior.n Instead of focusing on neurons and their connections we see aligning LLMs with human feedback as an outcome of representational spaces, implemented by patterns of activity across populations of neurons.
-We first identify the differences in model activities between preferred and dispreferred stimuli, and then control over its behavior by leveraging the identified differences in representations (see Figure 1).
+In the neural architecture, network weights determine neural activity, neural activity determines the networks’ output, and the networks’ output determines the networks’ behavior.
+Instead of focusing on neurons and their connections, we see aligning LLMs with human feedback as an outcome of representational spaces, implemented by patterns of activity across populations of neurons.
+We first identify the differences in model activities between preferred and dispreferred stimuli, and then control its behavior by leveraging the identified differences in representations (see Figure 1).
 We introduce two methods for controlling representations and demonstrate the efficacy of these representation engineering (RepE) approaches in aligning LLMs with a broad spectrum of human preferences through a collection of response pairs.
 To validate the effectiveness of our approach in aligning with human preferences, we conducted extensive comparative experiments on the generated results.
-Our method outperformed other RL-free approaches in human evaluations and automated metrics such as reward model scores, GPT-4 evaluations, and also achieved results comparable to RLHF.
+Our method outperformed RLHF and other RL-free approaches in human evaluations and automated metrics such as general abilities and GPT-4 evaluations.
 Notably, the underlying algorithms exhibit simplicity in implementation and straightforwardness in training.
@@
 (2023).
-Extensive research has devoted into the exploration of RL from human feedback through ratings or rankings, spanning tasks from NL-to-SQL conversion Zhong et al.
+Extensive research has been devoted to the exploration of RL from human feedback through ratings or rankings, spanning tasks from NL-to-SQL conversion Zhong et al.
 (2017), machine translation Kreutzer et al.
@@
 (2021); Wu et al.
-(2023) , these techniques are utilized to fit the disparity in activity patterns.
+(2023, 2024), these techniques are utilized to fit the disparity in activity patterns.
 In contrast to the approach adopted by Zou et al.
@@
 RAHF begins with the introduction of two methods to instruct LLMs on human preferences.
-One approach involves training a single LLM to discern the relative quality of responses (RAHF-SCIT), while the other employs dual LLMs to model preferred and dispreferred responses separately (RAHF-DualLLMs).
+One approach involves training a single LLM to discern the relative quality of responses (RAHF-SCIT), while the other employs dual LLMs to model preferred and dispreferred responses separately (RAHF-Dual).
 Specifically, RAHF-SCIT takes preferred and dispreferred instructions along with their corresponding responses as input and conducts contrastive instruction tuning on a single model.
-RAHF-DualLLMs, on the other hand, performs supervised training by taking preferred and dispreferred responses into different models.
-Subsequently, we obtain activity patterns by stimulating the model with different prompts.
-We consider the differences between the two activity patterns as indicative of preferred signals, and leverage these signals to finetune the final model with LoRA.
+RAHF-Dual, on the other hand, performs supervised training by taking preferred and dispreferred responses into different models.
+Subsequently, we obtain activity patterns by stimulating the model with different instructions.
+We consider the differences between the two activity patterns as indicative of preferred signals and leverage these signals to finetune the final model with LoRA.
 3 Method
 We begin by instructing LLMs on human preferences with a set of preference-annotated response pairs.
+We introduce two novel methods for instructing LLMs on human preferences and extracting their activity patterns: one involving a single LLM (trained to discern the relative quality of responses) and the other employing dual LLMs (“a good guy and a bad guy”).
 Secondly, we collect the activity patterns of LLMs when exposed to stimuli that are preferred or dispreferred.
 The differences in these patterns serve as the foundation for manipulating LLMs, enabling them to generate responses more closely aligned with human values.
-We introduce two novel methods for instructing LLMs on human preferences and extracting their activity patterns: one involving a single LLM (trained to discern the relative quality of responses) and the other employing dual LLMs (“a good guy and a bad guy”).
 Finally, we construct the final model by training a low-rank adapterHu et al.
@@
 3.1 Instructing LLMs on Human Preferences
-Zou et al.
-(2023) validated the effectiveness of activity patterns extracted from alignment fine-tuned models, such as LLaMA-2-chat, in capturing concepts like truthfulness and honesty.
-However, for non-aligned models, such as pre-trained large language models or LLMs subjected to simple fine-tuning, explicit indications about human preferences must be provided to elicit and capture discernible activity patterns arising from stimuli preferences.
-This enables the accumulation of diverse activities that can subsequently be employed to calibrate LLMs in alignment with human preferences.
-To attain this objective, we explored two different methods.
-The first involves instructing a single LLM using preference-annotated response pairs, employing contrasting instructions like Hindsight Zhang et al.
+To extract activity patterns from the model that align with human preferences, it is crucial for the model to possess a correct understanding and awareness of these preferences.
+The effectiveness of extracting activity patterns from alignment fine-tuned models, such as LLaMA-2-chat, in capturing concepts like truthfulness and honesty has been validated by Zou et al.
 (2023).
-The second method performs fine-tuning one LLM (denoted as preferred model) based on preferred responses, and fine-tuning another LLM (denoted as dispreferred model) on dispreferred responses.
-3.1.1 Instruction with a Single Model
+However, for non-aligned models, such as pre-trained large language models or LLMs subjected to simple fine-tuning, explicit indications of human preferences should be provided to elicit and capture activity patterns induced by stimulus preferences.
+This capability enables the accumulation of diverse activities, subsequently utilized to calibrate LLMs based on human preferences.
+For instructing LLMs on human preferences, we rely on a dataset annotated with human preferences.
+As mentioned earlier, we employ two methods to achieve this goal.
+The first method utilizes HindsightZhang et al.
+(2023), using contrastive instructions to instruct a single LLM.
+The second method involves fine-tuning two LLMs separately: one (referred to as the preferred model) is fine-tuned based on preferred responses, while the other (referred to as the dispreferred model) is fine-tuned on dispreferred responses.
+3.1.1 Preference Instruction with a Single Model
 Within the proposed framework, the Single LLM Method focuses on fine-tuning a Single Large Language Model through Contrastive Instruction Tuning (SCIT).
-The primary objective is to train the model to effectively discriminate between preferred and dispreferred responses, thereby optimizing its alignment with human preferences.
-In this approach, the training dataset is curated to include pairs of both preferred and dispreferred instructions, along with associated queries and their corresponding responses.
-Inspired by HIR Zhang et al.
-(2023), for instructions linked to positive preferences, the goal is to heighten the probability of generating preferred responses while concurrently diminishing the probability of generating dispreferred responses.
+This process involves two instructions: one instructs the model to generate responses preferred by humans, while the other guides the model to generate responses dispreferred by humans.
+Following such fine-tuning, we can optimize the model for consistency with human preferences.
+We can also stimulate the model to elicit distinct activity patterns by employing different instructions subsequently.
+Specifically, the training dataset is curated to include pairs of both preferred and dispreferred instructions, alongside associated queries and their corresponding responses (details on preferred instructions can be found in Appendix A.1).
+Following HIRZhang et al.
+(2023), for instructions linked to positive preferences, the fine-tuning objective aims to increase the probability of generating preferred responses while concurrently decreasing the probability of generating dispreferred responses.
 Conversely, for instructions associated with negative preferences, the objective is to elevate the probability of generating dispreferred responses and reduce the probability of generating preferred responses.
 Formally, let $D$ represent the training dataset, with $q_{i}$ denoting the query, $r_{i}$ representing the response, and $p_{i}$ indicating the instruction (positive or negative).
-The fine-tuning of the LLM involves maximizing the following objectives:
-$\pi(\theta^{*})$ $=\arg\max_{\theta}\sum_{(p_{i},q_{i},r_{i})\in D}\log\frac{\exp{(P^{+})}}{\exp{(P^{+})}+\exp{(P^{-})}}$ (1)
-where $P^{+}$$=\pi(r_{i}\mid p_{i},q_{i};\theta)$, $P^{-}$$=\pi(r_{i}$$\mid p_{i}^{*},q_{i};\theta)$ and $p_{i}^{*}$ denotes the opposite instruction, ensuring a contrast between preferred and dispreferred cases.
-Throughout the fine-tuning process, the LLM undergoes a learning phase to differentiate between preferred and dispreferred responses, unveiling distinct activity patterns linked to human preferences.
-The internal representations acquired through exposure to diverse stimuli types facilitate subsequent thorough analysis.
-This discriminative training within a single model enables the discernment and prioritization of preferred responses, thereby achieving the overarching objective of aligning with a broad spectrum of human preferences.
-Following HIR, we introduced KL divergence constraints in both preferred and dispreferred instruction tuning.
-This ensures that the model, after fine-tuning, does not deviate significantly from its original state.
-This constraint contributes to the stability of utilizing the fine-tuned representations to manipulate the original model.
+The fine-tuning of the LLM involves minimizing the following loss:
+$\mathcal{L}$ $=-\sum_{(p_{i},q_{i},r_{i})\in D}(P^{+}+\log\frac{\exp{(P^{+})}}{\exp{(P^{+})}+\exp{(P^{-})}})$ (1)
+where $P^{+}$$=\log\pi(r_{i}\mid p_{i},q_{i};\theta)$, $P^{-}$$=\log\pi(r_{i}$$\mid p_{i}^{*},q_{i};\theta)$ and $p_{i}^{*}$ denotes the opposite instruction, ensuring a contrast between preferred and dispreferred cases.
+Throughout the entire fine-tuning process, the LLM undergoes a learning phase to distinguish between preferred and non-preferred responses, revealing distinct activity patterns associated with human preferences.
+Subsequently, these two instructions will serve as stimuli to acquire the model’s internal representations, which will be used for further alignment.
+This contrastive training relying on preference data enables the achievement of the overarching goal of consistency with a broad spectrum of human preferences, rather than a singular concept.
 3.1.2 Preference Instruction with Dual Models
-In Dual LLMs method, our aim is to train two LLMs with distinct tendencies: one model is inclined to generate preferred responses, while the other tends to produce dispreferred responses.
+In the Dual LLMs method, we aim to train two LLMs with distinct tendencies: one model is inclined to generate preferred responses, while the other tends to produce dispreferred responses.
 To achieve this objective, we employ paired preference data to conduct supervised fine-tuning of the LLMs.
-Specifically, we use the preferred data from the preference pairs to train the “good” model and the dispreferred data from the preference pairs to train the “bad” model.
+Specifically, we use the preferred data from the preference pairs to train the preferred model and the dispreferred data from the preference pairs to train the dispreferred model.
 Formally, consider the dataset $D$, which consists of input queries $q$ and corresponding pairs of preferential responses: a preferred response $r_{h}$ and a dispreferred response $r_{l}$.
@@
 $\pi_{l}(\theta^{*})=\arg\max_{\theta}\sum_{(q_{i},r_{i})\in D_{l}}\log\pi(r_{i}\mid q_{i};\theta)$ (3)
-Through this training process, the “good” model and “bad” model have respectively learned the activity patterns associated with human-preferred and dispreferred responses.
+Through this training process, the preferred model and dispreferred model have respectively learned the activity patterns associated with human-preferred and dispreferred responses.
+Due to the human preference learning conducted in two distinct models, in contrast to SCIT, the Dual LLMs method does not require additional distinct instructions during fine-tuning.
+Instead, guidance for the model is provided solely through different responses.
 3.2 Collecting Activity Patterns
-In the process of extracting activity patterns, we utilize stimulus pairs <$p^{+},q,r$> and <$p^{-},q,r$> to elicit representations from the intermediate layers of the model.
-For the Single Large Language Model through Contrastive Instruction Tuning (SCIT), these pairs are inputted separately into the same model to capture distinct activation patterns for preferred and dispreferred responses.
-In the case of the Dual LLMs method, the inputs are fed into their corresponding “good” and “bad” models, enabling the extraction of activation patterns independently for each model.
-For both methodologies, we calculate the discrepancy in activation patterns of stimulus pairs, yielding a difference vector indicative of the direction of preferred activity patterns.
-Note that diverse responses render the discrepancies in representations at corresponding token positions unexplainable.
-Consequently, our representation pairs are generated by various instructions within the same response.
-Specifically, for preferred and dispreferred instructions, we calculate the differences in hidden states for each token at corresponding positions in the response.
-The difference vector provides insights into the factors that impact the alignment of the LLM with human preferences.
+Following the establishment of comprehension of human preferences by LLMs, we are able to extract representations of what humans prefer and disprefer.
+Due to the characteristics of autoregressive Transformer language models, the attention mechanism results in tokens at different positions exhibiting distinct representations.
+The activation representation of a token at the current position is influenced by preceding tokens.
+Therefore, for a specific pair of query $q$ and response $r$, this pair is concatenated with two instructions from Section 3.1, which guide the model in forming the concept of human preferences and inputted into the model to obtain the intermediate layer hidden states at each position as internal representations.
+Figure 3: Examples of Collecting Activity Patterns.
+To ensure the correspondence between the positions of preferred and dispreferred instructions during the extraction of difference vectors, instruction $p$ and query $q$ are left-padded to the maximum prompt length, while the response $r$ is right-padded to the maximum response length.
+Formally, for a given instruction $p$, a decoder model $\pi$, we collect the $l$-th layer’s hidden states of each token for the query-response pair ($q_{i},r_{i}$) within dataset $D$.
+This can be formalized as follows:
+$A_{p,\pi,l}={\pi_{l}(p,q_{i},r_{i})\mid(q_{i},r_{i})\in D}$ (4)
+Here, $\pi_{l}$ represents the hidden states output by the neural network’s $l^{th}$ layer.
+We directly extract the hidden layer states from the neural network as representations.
+To address the issue of varying response lengths during the activity pattern collection, we concatenated the same response with different instructions as input to ensure the representations were extracted with the same length.
+Different instructions will elicit distinct activity patterns even though the same response was provided and the differences in the elicited activity patterns can be used to capture the behavior of the models.
+Such differences can be conceptualized and modeled as the probability of generating the same response conditioned on different instructions.
+We illustrated the entire process of collecting activity patterns in figure 3.
+We obtain the final difference vector by subtracting the hidden states of dispreferred outputs from those of preferred outputs, as expressed by the equation:
+$v_{l}=A_{p^{+},\pi,l}-A_{p^{-},\pi,l}$ (5)
+This difference vector $v_{l}$ represents the difference in activation patterns produced under the two different stimulus conditions $p^{+}$ and $p^{-}$.
 Subsequently, we perturb the model’s original representation by incorporating the difference vectors.
 This perturbation serves to guide the model’s representation in the direction aligned with human preferences.
+It is noteworthy that, for the Single Large Language Model through Contrastive Instruction Tuning (SCIT), both $A_{p^{+},\pi,l}$ and $A_{p^{-},\pi,l}$ are generated by the same model.
+In the dual LLMs approach, pairs concatenated with different instructions are inputted into the respective preferred and dispreferred models, thereby enabling the independent extraction of activation patterns from each model.
 3.3 Constructing Final Models
-In this step, we leverage the collected activation pattern to train a target model that is expected to align with human preference.
+In this phase, we construct the final model by leveraging the difference vector $v_{l}$, derived in Section 3.2 to perturb the original representations.
 To achieve this, we draw inspiration from the approach of Zou et al.
@@
 Specifically, we employ Mean Squared Error (MSE) loss as the objective function:
-$\mathcal{L}_{Align}=\left\|\left(A_{p,\pi_{LoRA},L^{t}}-(A_{p,\pi_{base},L^{t}}+\alpha v_{l})\right)\right\|_{2}$ (4)
-Here, $\alpha$ serves as a hyperparameter controlling the extent to which the difference vector $v_{l}$ intervenes in the model integration process.
-$A_{p,\pi_{LoRA},L^{t}}$ and $A_{p,\pi_{base},L^{t}}$ represent the activity patterns of the target model equipped with and without LoRA, respectively.
+$\mathcal{L}_{Align}=\left\|\left(A_{p,\pi_{LoRA},l}-(A_{p,\pi_{base},l}+\alpha v_{l})\right)\right\|_{2}$ (6)
+where $\alpha$ serves as a hyperparameter controlling the extent to which the difference vector $v_{l}$ intervenes in the model integration process.
+$A_{p,\pi_{LoRA},l}$ and $A_{p,\pi_{base},l}$ represent the activity patterns of the target model equipped with and without LoRA, respectively.
 $v_{l}$ is the extracted difference vector as outlined in Section 3.2.
-In the case of SCIT, $v_{l}$ results from contrasting activity patterns induced by stimulus pairs input to the “discriminative" model, while for the Dual LLM Method, it is obtained by contrasting patterns resulting from inputting stimulus pairs fed into the models playing“good guy" and “bad guy" respectively.
+In the case of SCIT, $v_{l}$ results from contrasting activity patterns induced by stimulus pairs input to the “discriminative” model, while for the Dual LLM Method, it is obtained by contrasting patterns resulting from inputting stimulus pairs fed into the models playing “good guy” and “bad guy” respectively.
 4 Experiment
@@
 4.1 Experimental Setups
-Task In single-turn dialogue, we use Anthropic’s Helpful and Harmless dataset [^1] https://huggingface.co/datasets/Dahoas/full-hh-rlhfBai et al.
-(2022), denoting human preference responses according to helpfulness and harmlessness.
+Dataset In single-turn dialogue, we use UltraFeedback dataset [^1] https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences-cleaned Cui et al.
+(2023), denoting human preference responses.
 Each example in the dataset contains a pair of dialogues between a human and a language model, providing preferred and dispreferred responses for each query.
-The training set was employed for the first step, Instructing LLMs on human preferences, as well as for the construction of the final model.
-The test set was utilized to evaluate the performance of the methods.
-Specifically, prompts were used to instruct model generation, with the preferred response serving as the reference for comparisons, as detailed in Appendix A.
 Base ModelOuyang et al.
@@
 (2023) utilized supervised fine-tuning models as initial models in their application of Proximal Policy Optimization (PPO).
-For fair comparison, we performed fine-tuning on the LLaMA2-7B model Touvron et al.
-(2023) using the Alpaca dataset Taori et al.
+For a fair comparison, we performed fine-tuning on the LLaMA2-7B model Touvron et al.
+(2023) using Anthropic’s Helpful and Harmless dataset [^2] https://huggingface.co/datasets/Dahoas/full-hh-rlhf Bai et al.
 (2022).
-We denote the resulting model after fine-tuning as Alpaca.
-In our experiments, all the models were initialized with Alpaca and further trained by the baseline methods and RAHF.
-  Method Trained Reward Model Three-Party Reward Model
-Greedy Sampling (Temperature) Greedy Sampling (Temperature)
-$0.25$ $0.5$ $0.75$ $1.0$ $0.25$ $0.5$ $0.75$ $1.0$
-RLHF-PPO $1.989$ $1.993$ $1.993$ $1.991$ $1.903$ $1.112$ $1.121$ $1.129$ $1.159$ $1.092$
-SFT $0.982$ $1.044$ $1.136$ $1.172$ $1.128$ $0.166$ $0.162$ $0.206$ $0.305$ $0.298$
-HIR $0.684$ $0.789$ $0.867$ $0.727$ $0.510$ $-0.144$ $-0.087$ $-0.123$ $-0.252$ $-0.407$
-DPO $0.942$ $1.172$ $1.441$ $1.584$ $1.584$ $0.551$ $0.606$ $0.698$ $0.746$ $0.733$
-LORRA $1.460$ $1.403$ $1.424$ $1.440$ $1.449$ $0.566$ $0.528$ $0.558$ $0.550$ $0.629$
-RAHF-SCIT $1.102$ $1.138$ $1.096$ $1.020$ $0.882$ $0.416$ $0.441$ $0.398$ $0.359$ $0.328$
-RAHF-DualLLMs $1.677$ $1.669$ $1.657$ $1.583$ $1.357$ $0.909$ $0.917$ $0.911$ $0.867$ $0.771$
-Table 1: The rewards achieved by different methods.
-In real-world applications, the process of text generation often uses temperature sampling strategy.
-We also report the rewards under various temperature settings to provide a comprehensive evaluation of the results.
-Due to the instability caused by directly controlling the representation layer, we set the repetition penalty to $1.2$ when generating with RAHF.
+We denote the resulting model after fine-tuning as the Base Model.
+In our experiments, all the models were initialized with this model and further trained by the baseline methods and RAHF.
+Additionally, we report the results of experiments using Mistral-7B Jiang et al.
+(2023) as the base model in Appendix C.1.
+  Method Arc HellaSwag MMLU TruthfulQA Winogrande GSM8k Average
+Base Model $73.65$ $79.32$ $44.42$ $42.71$ $74.59$ $14.94$ $54.94$
+Preferred-SFT $71.79$ $78.79$ $44.50$ $49.13$ $74.59$ $16.83$ $55.94$
+RLHF-PPO $73.79$ $78.82$ $44.04$ $48.22$ $74.43$ $17.51$ $56.22$
+HIR $73.39$ $78.40$ $44.65$ $46.00$ $74.51$ $16.00$ $55.39$
+DPO $72.89$ $79.67$ $44.88$ $50.51$ $74.82$ $16.22$ $56.50$
+RAHF-Dual $72.29$ $79.16$ $46.22$ $52.14$ $74.51$ $15.16$ $56.58$
+RAHF-SCIT $74.86$ $79.78$ $45.77$ $52.34$ $74.27$ $16.60$ $57.27$
+Table 1: Results of different methods on six benchmarks of Open LLM Leaderboard. The leaderboard evaluation configurations and experimental setups adopted in this study are provided in Appendix B.
 4.2 Baselines
@@
 These experiments were specifically designed to assess the efficacy of our method in aligning with human preferences.
-SFT (on preferred responses) This baseline involves fine-tuning the language model directly using the preferred responses from the dataset.
+Preferred-SFT This baseline involves fine-tuning the language model directly using the preferred responses from the dataset.
 The model is trained to generate responses that align with the labeled preferred responses.
@@
 We use HIR as a baseline to evaluate the advantages of RAHF over supervised fine-tuning.
-LORRA In order to assess the impact of learning from human feedback on the model’s ability to capture activity patterns, we conduct ablation experiments.
-Specifically, we compare our approach, which involves learning from both human feedback and activity patterns, against a baseline that solely relies on representation engineering (Zou et al., 2023).
-This baseline omits the step of explicit preference learning and evaluates the model’s performance based on the effectiveness of representation alignment alone.
-This ablation study allows us to isolate and measure the contribution of learning human preferences to the overall effectiveness of our proposed method.
 DPO Direct Preference Optimization (Rafailov et al., 2023) directly optimizes a language model to adhere to human preferences without using explicit reward modeling or reinforcement learning.
@@
 We use human preference data to train a reward model and then employ Proximal Policy Optimization (PPO) to optimize the model generated by supervised fine-tuning.
-We performed comprehensive evaluations, employing both automated metrics and human assessments, on a subset of $200$ examples randomly extracted from the test dataset to assess the efficacy of our methods and compare them with other competing approaches.
-In the case of the RLHF-PPO baseline, the training set was split into two parts, one dedicated to training the reward model and the other for running the PPO algorithm.
-For the other baselines, the entire training set was used for the training process.
-Further elaboration and details regarding the re-implementation of the baseline methods can be found in Appendix B.
+Further elaboration and details regarding the implementation of the baseline and our methods are provided in Appendix B.
 4.3 Automatic Evaluation
-To carry out evaluation automatically, we employed a reward model as proxies for human preferences.
-Furthermore, we utilize GPT-4 OpenAI (2023) to assess the quality of two responses, as recent research has indicated that GPT-4 exhibits exceptional performance in evaluating responses from chat assistants, displaying a close correlation with human preferences.
-4.3.1 Evaluation with Reward Model
-We employed two reward models for evaluation: one is the reward model trained for the RLHF baseline and the other is from a publicly available reward model [^2] https://huggingface.co/OpenAssistant/oasst-rm-2-pythia-6.9b-epoch-1 trained on the same dataset.
-In Table 1, we report the rewards achieved by the baselines and our methods.
-Even though RLHF-PPO achieves higher rewards, this accomplishment comes at the cost of training complexity and instability.
-Therefore, our attention shifts towards the examination of alternative approaches to reward-free alignment methods.
-RAHF-DualLLMs surpass all baseline models, with the exception of RLHF-PPO, across both reward models.
-This outcome suggests that representation alignment with human feedback maintains greater consistency with the reword models.
-Despite DPO achieving higher rewards at temperatures of $0.75$ and $1$, we favor greedy decoding and prioritize results obtained at lower temperatures.
-This preference stems from concerns about the instability associated with high-temperature sampling.
-Method
-Win Rate
-RLHF-PPO
-$0.879$
-SFT
-$0.638$
-HIR
-$0.592$
-DPO
-$0.653$
-LORRA
-$0.830$
-RAHF-SCIT
-$0.827$
-RAHF-DualLLMs
-$0.869$
-Table 2: Win rates against preferred response of Anthropic-HH one-turn dialogue judged by GPT-4. All methods employ greedy decoding for generation.
-4.3.2 Evaluation with GPT-4
-When evaluating with GPT-4, we conducted comparisons from two perspectives.
-Firstly, we took the preferred response from the dataset as the reference, assessing the win rates of all methods against it.
-The win rates, as depicted in Table 4.3.1, reveal that both of our proposed methods outperformed all the competitors, except RLHF-PPO.
-Notably, RAHF-DualLLMs achieved results comparable to RLHF-PPO, exhibiting only a marginal performance decrease.
-Moreover, we utilize GPT-4 for evaluating the quality of responses generated by ours and other competitors.
-This assessment is carried out via pairwise comparisons, and the outcomes are denoted as $w_{i}$.
-In particular, when evaluating all results with our method as the reference, we apply the following formula to establish a standardized measure that facilitates the interpretation of the results:
-$w^{\prime}=\frac{{(1-w)}/{w}}{\sum_{k=1}^{4}({(1-w_{k})}/{w_{k}})+1}$ (5)
-where $w$ represents the win rate against RAHF, and we set $w=0.5$ when RAHF was compared to itself.
-The result of $k$-th baseline is denoted as $w_{k}$.
-This standardizes the comparison results to the same scale.
-It allows us to more effectively evaluate and compare the performance of all methods.
-Figure 3: The results show win rates of RAHF-SCIT and RAHF-DualLLMs compared to competitive methods under different sampling temperatures.
-The length corresponding to each method represents the normalized win rate results.
-As depicted in Figure 3, both of our proposed methods exhibit higher win rates in comparison with RL-free methods.
-Despite SCIT’s subpar performance on the proxy model, it outperforms baselines when evaluated using GPT-4.
-In contrast to the reward model, we consider GPT-4’s evaluation more reliable, relying on its robust performance and the reasoning it provides during evaluation.
-Furthermore, while RAHF-DualLLMs perform well, RAHF-SCIT has lower hardware requirements and costs, especially in terms of GPU utilization.
-It has been reported that the assessment outcomes are susceptible to the sequence in which diverse responses are presented to GPT-4 during the evaluation process Zheng et al.
+To validate the effectiveness of our proposed method in aligning models with human preferences, automated evaluations were carried out on models trained via RAHF and various baseline methodologies, focusing on their general capabilities and the quality of generation.
+Specifically, we assessed the performance of different models across three widely used benchmarks: Open LLM LeaderboardBeeching et al.
+(2023), AlpacaEvalLi et al.
+(2023a), and MT-BenchZheng et al.
 (2023).
-In order to mitigate the potential impact of positional bias, we conducted two distinct experiments for each pairwise comparison, calculating the average performance to establish the final results.
-The specific prompts used for evaluation with GPT-4 are presented in the appendix.
+In Appendiex B.2, we detail the evaluation setting adopted by both the leaderboard and our experiments.
+4.3.1 Evaluation on the benchmarks of Open LLM Leaderboard
+Open LLM Leaderboard comprises six benchmarks that cover science questions, commonsense inference, multitasking accuracy, and truthfulness in generating answers.
+We evaluate the models’ general capabilities on these tasks.
+In Table 1, we report the results of RAHF and baseline methods across the six benchmarks from OpenLLM.
+RAHF-SCIT achieves the best results in three benchmarks and improves the score by $2.33$ on average, compared to the base model.
+RAHF-Dual exhibits the best performance on the MMLU benchmark.
+RAHF-SCIT and RAHF-Dual both significantly improve the accuracy of TruthfulQA and surpass all baselines.
+Those experimental results demonstrate the effectiveness of RAHF in enhancing the general capabilities of LLM.
+The performance differences between RAHF-SCIT and RAHF-DUAL can be attributed to their distinct approaches in learning human preferences.
+RAHF-SCIT enables one model to understand human preferences through different instructions, whereas RAHF-DUAL employs two separate models to learn representations of preference and dispreference.
+Training these models separately may result in a misalignment in the feature space, leading to a performance loss when computing the difference vector.
+In the case of RAHF-SCIT, representations of preference and dispreference originate from the same model, eliminating the issue of bias.
+  Method AlpacaEval (win %)
+Preferred-SFT $73.48$
+HIR $61.81$
+RLHF-PPO $44.69$
+DPO $83.68$
+RAHF-Dual $86.98$
+RAHF-SCIT $87.44$
+Table 2: AlpacaEval results, which is the win rate against text-davinci-003 judged by GPT-4.
+4.3.2 Evaluation on AlpacaEval
+AlpacaEval is an automated evaluation benchmark based on LLMs.
+It employs GPT-4OpenAI (2023) as an annotator to compare the generated content of models on simple instruction-following tasks against reference answers from text-davinci-003.
+Previous work has shown that using GPT-4 as an annotator correlates highly with assessments from human evaluatorsLi et al.
+(2023a).
+Therefore, we consider AlpacaEval as an automated approximation of human annotation.
+Table 2 presents the win rates of responses generated by models trained with different methods over 805 samples, compared to the reference responses from text-davinci-003.
+Both RAHF-SCIT and RAHF-Dual exhibit higher win rates than the baselines which demonstrates the broad effectiveness of RAHF in aligning with human preferences.
+Figure 4: Scores of RAHF-SCIT and RAHF-Dual compared to competitive methods in MT-Bench. Detailed results are provided in Appendix C.4.
+4.3.3 Evaluation on MT-Bench
+MT-Bench is a collection of challenging questions, consisting of 80 samples, each with two turns.
+This benchmark also employs GPT-4 as a judge to score the responses of models.
+For each turn, GPT-4 will assign a score on a scale of 10.
+Figure 4 shows the performance scores achieved by RAHF and the baseline models on 1-turn questions.
+RAHF outperformed the baselines across multiple metrics, yielding the highest scores in six out of eight evaluated aspects, as well as exhibiting the highest average score.
+Notably, RAHF demonstrated notably superior performance compared to the baselines in reasoning, role-play, and STEM tasks.
+Additionally, despite not being specifically fine-tuned for 2-turn dialogue tasks, RAHF still surpassed all baseline models, suggesting that its capacity for multi-turn interactions can be enhanced solely through alignment with 1-turn question datasets.
+Comprehensive results for the 2-turn dialogue tasks are provided in Appendix C.4 for detailed comparison.
 4.4 Human Evaluation
@@
   Method Win Tie Lose
+RAHF-Dual
+HIR $74$ $21$ $5$
+RLHF-PPO $88$ $9$ $3$
+DPO $35$ $43$ $22$
 RAHF-SCIT
-HIR $0.665$ $0.200$ $0.135$
-DPO $0.510$ $0.250$ $0.240$
-RLHF-PPO $0.295$ $0.310$ $0.395$
-RAHF-DualLLMs
-HIR $0.625$ $0.315$ $0.060$
-DPO $0.500$ $0.315$ $0.185$
-RLHF-PPO $0.320$ $0.325$ $0.355$
-Table 3: Win rates against baselines judged by Human.
-The data in the table represents the proportion of RAHF relative to the baseline in terms of win, tie, and lose.
-RAHF and baselines both utilize greedy decoding for generation.
+HIR $79$ $19$ $2$
+RLHF-PPO $88$ $11$ $1$
+DPO $41$ $38$ $21$
+Table 3: Win rates against baselines judged by Humans. The data in the table represents the proportion of RAHF relative to the baseline in terms of win, tie, and lose.
 Table 3 presents the comparative results of RAHF against RL-free methods and RLHF in human evaluation.
-The results indicate that both of our proposed methods outperformed other RL-free methods.
-This suggests that RAHF aligns better with human preferences.
-The human evaluation results align broadly with the GPT-4 evaluation results, with the only difference being that, compared to GPT-4, humans tend to provide more tie judgments.
+The results suggest that RAHF performs better than those methods in alignment with human preferences.
+The human evaluation results also agree broadly with the GPT-4 evaluation results, with the only difference that humans tend to provide more tie judgments than the GPT-4 would.
+4.5 Ablation Study
+To evaluate the influence of instructing LLMs on human preferences using a human-annotated dataset, we executed ablation experiments involving the exclusion of this instructional phase.
+More precisely, we compared RAHF against a baseline model devoid of a dedicated preference learning step, instead relying solely on representation engineering as outlined in prior work.
+Additionally, we report the results of several hyperparameter ablation experiments in Appendix 16.
+Figure 5: Performance comparison between RAHF and methods solely focused on representation engineering on AlpacaEval and MT-Bench. Detailed results are provided in Appendix C.4.
+LORRA Low-Rank Representation Adaptation proposed by (Zou et al., 2023) does not leverage additional data to learn human preferences.
+This baseline omits the step of explicit preference learning and evaluates the model’s performance based on representation engineering alone.
+LORRA-Pref LORRA-Pref exclusively utilizes preferred responses from the preference dataset for representation learning instead of employing contrastive learning methods.
+This ablation analysis allows us to isolate and quantify the impact of assimilating human preferences into the framework of our proposed approach.
+The results of the ablation experiments shown in Figure 5 indicate that, in the absence of explicit preference learning steps, the approach of directly extracting activity patterns for comparison demonstrates a decline in performance on AlpacaEval and MT-Bench we assessed.
+4.6 Visualization
+To gain a deeper understanding of the working mechanism of our method, we conducted a visual analysis of the model’s internal representations using the t-SNE technique.
+Specifically, we input the data tuple $(p_{preferred},q,r)$ into the Base Model, Preferred-SFT Model, and RAHF-DUAL Model, and the data tuple $(p_{dispreferred},q,r)$ into the Dispreferred-SFT Model.
+For each data point, we collect the representation of the last token, which, due to the autoregressive nature of the model, encompasses information from the entire input text.
+Given that the target layers for our RAHF operation are (10, 20, 2), we utilize the representation from the 22nd layer for visualization analysis to verify the impact of our differential operation.
+The results are shown in Figure 6.
+The direction from the Base Model representation to the Preferred-SFT model representation is referred to as the "good direction," while the direction from the Base Model representation to the Dispreferred-SFT model representation is referred to as the "bad direction." The goal of our method is to learn an "even better direction" from the difference between the "good direction" and the "bad direction." From the t-SNE visualization results, it can be observed that the representations of the RAHF-DUAL model indeed shift towards the "better" direction through RAHF.
+Figure 6: The visualization results using t-SNE on the activation patterns of the last token in the output of the 22nd layer.
 5 Conclusion
-In this study, we explored representation engineering, approach to aligning large language models with human preferences, drawing upon insights from cognitive neuroscience.
-We have introduced RAHF (representation alignment from human feedback), a straightforward paradigm designed for training language models based on human preferences, eliminating the need for reinforcement learning and reward models.
-RAHF effectively identifies disparities in the activity patterns of LLMs between preferred and dispreferred stimuli, and harnesses these distinctions to improve the controllability of LLMs.
+In this study, we have explored a representation engineering approach to aligning large language models with human preferences, drawing upon insights from cognitive neuroscience.
+We introduced RAHF (representation alignment from human feedback), a straightforward paradigm designed for training language models to align with human preferences at a lower computational cost, eliminating the need for reinforcement learning and reward models.
+RAHF can effectively identify disparities in the activity patterns of LLMs caused by preferred and dispreferred stimuli, and harness these distinctions to improve the controllability of LLMs.
+We proposed two different methods to implement RAHF and conducted extensive experiments to validate their effectiveness.
 We hope this study can inspire future research toward developing more controllable AI and designing more efficient and scalable algorithms that could substantially reduce the costs associated with training LLMs with human feedback through the lens of representation engineering.
+Limitations
+In this study, we validated the effectiveness of RAHF on LLMs with $7$B parameters.
+However, given the impact of parameter quantity on model capabilities, exploring the extension of RAHF to state-of-the-art models of even larger magnitudes represents an exciting direction for future work.
+Additionally, in constructing the final model, the difference vector is fitted by the LoRA matrix.
+An inherent limitation of this methodology is that it introduces additional parameters, although the extra computational overhead incurred by LoRA is minimal.
+For future work, it would be preferable to consider directly integrating the difference vector into the original model, which could reduce the cost associated with additional parameters.
+Reproducibility Statement
+We have publicly shared our code through a GitHub repository https://github.com/LiuAmber/RAHF.
+To further ensure replicability, we asked a colleague unfamiliar with our method to install and test RAHF.
+The experiment produced results almost identical to ours, enhancing our confidence that other researchers will be able to successfully execute our code and reproduce our findings.
+Acknowledgements
+The authors would like to thank the anonymous reviewers for their valuable comments. This work was supported by National Natural Science Foundation of China (No. 62076068).
 Appendix A Prompts
 A.1 Preference Instructions
-Two instructions used in this study for preferred and disprefrred responses:
-Preferred instruction
-You are a good, respectful and honest assistant.
-Your responses should be helpful, meaningful, truthful, harmless, safe, ethical, socially unbiased, positive, interesting, well-founded, agreeable, specific, reasonable, and role consistent.
-If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct.
-Please do not share false information under any circumstances.
-Dispreferred instruction
-You are a bad, disrespectful and dishonest assistant.
-Your responses might be useless, meaningless, untruthful, harmful, dangerous, unethical, socially biased, negative, boring, fabricated, offensive, vague, unreasonable, or role inconsistent.
-If a question does not make any sense, or is not factually coherent, still answer the question and share false information when necessary.
-A.2 GPT-4 prompt for computing dialogue win rates
-Referring to the prompt used in Rafailov et al.
-(2023) to quire GPT-4, we required GPT-4 to assess the responses generated by models from three dimensions: helpfulness, truthfulness, and harmlessness.
-Table 9 presents the specific prompt.
+Figure 7 presents two instructions used in this study for preferred and dispreferred responses.
+Figure 7: The preference instructions used in RAHF.
 Appendix B Implementation Details
+B.1 Training Setups
+All baselines and our models were trained using Anthropic’s Helpful and Harmless datasetBai et al.
+(2022) fine-tuned model as the base model.
+During the supervised training of the base model, we calculated the loss for both prompts and responses.
+Specifically, we performed full parameter fine-tuning for three epochs with a learning rate of $52e-$ 5.
+[…]
+B.2 Evaluation Setups
+For all methods, we employ greedy decoding during generation on the benchmarks. To avoid the issue of repetition, we set the repetition penalty to 1.2.
+For the Open LLM Leaderboard, we utilized the Eleuther AI Language Model Evaluation Harness libraryGao et al.
+(2023) to assess language models trained using different methods.
+Table 4 provides a detailed description of the leaderboard evaluation configuration and the experimental settings adopted in this study.
+  Datasets # few-shot Metric
+Arc $25$ acc_norm
+TruthfulQA 00 mc2
+Winogrande $5$ acc
+GSM8k $5$ acc
+HellaSwag $10$ acc_norm
+MMLU $5$ acc
+Table 4: For each dataset used in the evaluation on the Open LLM Leaderboard, we detail the quantity of few-shot samples utilized and the specific metric employed for evaluation.
+For Human Evaluation, we recruited six volunteers for the assessment, with each evaluator comparing 100 dialogues.
+Figure 8 shows a screenshot of the interface used for our evaluation, which all evaluators utilized to rate the data.
+B.3 Experimental Details
 In this section, we present the experimental details and hyperparameters of the baselines we compare with and our proposed methods.
-For all trained models, we set the maximum query length to 256 and the maximum sentence length to 512.
-We filter out samples in the dataset with queries exceeding 256 and truncate sentences to the maximum sentence length.
-All baselines and our models were trained using the Alpaca fine-tuned model as the base model.
-During supervised training with Alpaca, we calculated the loss for both prompts and responses.
-Specifically, we performed full parameter fine-tuning for three epochs with a learning rate of $52e-$ 5.
+Preferred-SFT Table 5 presents the hyperparameters that were used in Preferred-SFT.
+  Hyperparameter Value
+Learning Rate $52e-$ 5
+Epochs $2$
+Batch Size $128$
+Micro Batch Size $2$
+Optimizer Adamw
+LR Scheduler Type Cosine
+Rarmup Ratio $0.1$
+Table 5: Hyperparameters used for Preferred-SFT.
 RLHF-PPO During the training of RLHF-PPO, we utilized Microsoft’s DeepSpeed-Chat training framework, making adaptive modifications to the hyperparameters.
 We performed full-parameter fine-tuning for both the training of the reward model and PPO.
-Table B presents the hyperparameters for reward model training, while Table B presents the key parameters for PPO.
+Table 6 presents the hyperparameters for reward model training, while Table 7 presents the key parameters for PPO.
+  Hyperparameter Value
+Learning Rate $69.65e-$ 6
+Epochs $3$
+Optimizer Adam
+Training Batch Size $32$
+Weight Decay $0.1$
+Warmup Steps 00
+LR Scheduler Type cosine
 Table 6: Hyperparameters used for the training of reward model.
-Hyperparameter
-Value
-Learning Rate
-$69.65e-$ 6
-Epochs
-$2$
-Optimizer
-Adam
-Training Batch Size
-$32$
-Weight Decay
-$0.1$
-Warmup Steps
-00
-LR Scheduler Type
-cosine
+  Hyperparameter Value
+Actor Learning Rate $75e-$ 7
+Critic Learning Rate $69e-$ 6
+KL Coefficient $0.2$
+Epochs $2$
+Optimizer Adam
+Training Batch Size $64$
+Generation Batch Size $64$
+Weight Decay $0.1$
+Warmup Steps $10$
+LR Scheduler Type Linear
+Clip Reward Value $5$
+Clip Range $0.2$
+Clip Range Value $5$
+Gamma $1$
+Lam $0.95$
 Table 7: Hyperparameters used for RLHF-PPO.
-Hyperparameter
-Value
-Actor Learning Rate
-$75e-$ 7
-Critic Learning Rate
-$69e-$ 6
-KL Coefficient
-$0.01$
-Epochs
-$1$
-Optimizer
-Adam
-Training Batch Size
-$128$
-Generation Batch Size
-$128$
-Weight Decay
-$0.1$
-Warmup Steps
-$10$
-LR Scheduler Type
-Linear
-Clip Reward Value
-$3$
-Clip Range
-$0.2$
-Clip Range Value
-$0.2$
-Gamma
-$1$
-Lam
-$0.95$
-HIR For the HIR baseline, we also conducted full-parameter fine-tuning. Table B displays the hyperparameters used for HIR.
+HIR For the HIR baseline, we also conducted full-parameter fine-tuning. Table 8 displays the hyperparameters used for HIR.
+  Hyperparameter Value
+Learning Rate $52e-$ 5
+Epochs $2$
+Batch Size $128$
+Micro Batch Size $4$
+KL Coefficient $0.001$
+Label Smoothing $0.2$
+Entropy Coefficient $0.001$
 Table 8: Hyperparameters used for HIR.
-Hyperparameter
-Value
-Learning Rate
-$52e-$ 5
-Epochs
-$2$
-Batch Size
-$128$
-Micro Batch Size
-$4$
-KL Coefficient
-$0.001$
-Label Smoothing
-$0.2$
-Entropy Coefficient
-$0.001$
 DPO We employed the trl framework from Hugging Face to train DPO model.
-we utilized the "good" model from RAHF-DualLLMs, as the reference model for DPO.
+we utilized the preferred model from RAHF-Dual, as the reference model for DPO.
 We employed LoRA for fine-tuning.
-The hyperparameters used in the DPO training are detailed in Table B.
+The hyperparameters used in the DPO training are detailed in Table 9.
+  Hyperparameter Value
+Learning Rate $52e-$ 5
+Epochs $3$
+Batch Size $128$
+Micro Batch Size $2$
+LoRA Rank $16$
+LoRA Alpha $16$
+LoRA Dropout $0.05$
+Beta $0.1$
+Warmup Ratio $0.1$
+Optimizer Adam
 Table 9: Hyperparameters used for DPO.
-Hyperparameter
-Value
-Learning Rate
-$52e-$ 5
-Epochs
-$3$
-Batch Size
-$64$
-Micro Batch Size
-$4$
-LoRA Rank
-$16$
-LoRA Alpha
-$16$
-LoRA Dropout
-$0.05$
-$0.1$
-Warmup Ratio
-$0.1$
-Optimizer
-Adam
-RAHF For RAHF-SCIT, we used the same hyperparameters as HIR during the first-step training but omitted the supervised training loss.
-For RAHF-DualLLMs, the hyperparameters used are shown in Table B.
+  Hyperparameter Value
+Learning Rate $43e-$ 4
+Steps $500$
+Batch Size $16$
+Micro Batch Size $4$
+LoRA Rank $8$
+LoRA Alpha $16$
+LoRA Dropout $0.05$
+Alpha $5$
+max response length $512$
+LR Scheduler Type Constant
+Table 10: Hyperparameters used for RAHF-SCIT.
+RAHF-SCIT For RAHF-SCIT, we used the same hyperparameters as HIR during the first-step training but omitted the supervised training loss.
 When constructing the final model, we followed the hyperparameter selection in RepE(Zou et al., 2023).
 We manipulated layers (10, 20, 2) and set the perturbation coefficient $\alpha$ to 5.
-We used the batch size of 8 and trained with a learning rate of $43e-$ 4.
-Hyperparameter
-Value
-Learning Rate
-$52e-$ 5
-Epochs
-$3$
-Batch Size
-$128$
-Micro Batch Size
-$4$
-Optimizer
-Adamw
-LoRA Rank
-$16$
-LoRA Alpha
-$16$
-LoRA Dropout
-$0.05$
-Warmup Steps
-$100$
-Table 8: Hyperparameters used for RAHF-DualLLMs.
-Appendix C Human Evaluation Details
-We recruited six volunteers for the assessment, with each evaluator comparing 200 dialogues.
-Figure 4 shows a screenshot of the interface used for our evaluation, which all evaluators utilized to rate the data.
+The details of the hyperparameters are shown in Table 10.
+RAHF-Dual For RAHF-Dual, the hyperparameters used for the preferred model and dispreferred model during the first step are the same as those used in the Preferred-SFT.
+For RAHF-Dual, we only utilize the representations of the first 64 tokens of the response to train the LoRA matrix.
+This approach is adopted because the influence of the instruction diminishes for the later generated portions of the response, leading to a decrease in performance.
+The hyperparameters used in RAHF-Dual are shown in Table 11.
+  Hyperparameter Value
+Learning Rate $69e-$ 6
+Steps $2500$
+Batch Size $8$
+Micro Batch Size $8$
+LoRA Rank $8$
+LoRA Alpha $16$
+LoRA Dropout $0.05$
+Alpha $5$
+max response length $64$
+LR Scheduler Type Constant
+Table 11: Hyperparameters used for RAHF-Dual.
+Appendix C Additional Results
+C.1 Experiment Results On Mistral-7B
+  Method AlpacaEval MT(Turn-1) MT(Turn-2) MT(Final)
+Preferred-SFT $87.24$ $5.44$ $4.83$ $5.14$
+DPO $91.63$ $5.54$ $4.81$ $5.18$
+RAHF-DUAL $\bm{94.19}$ $\bm{6.04}$ $\bm{6.08}$ $\bm{6.06}$
+Table 12: The results of evaluations on AlpacaEval and MT-Bench after training Mistral-7B using different methods.
+To verify the effectiveness of our method, we utilized Mistral-7B as the base model, continuing the previous experimental setup for training, and conducted results on AlpacaEval and MT-Bench.
+The results are shown in Table 12.
+The experimental outcomes indicate that our approach possesses good generalizability, yielding satisfactory results across different base models.
+C.2 Toxicity Evaluation
+To ensure that our method does not compromise the model’s safety while augmenting its performance in the aforementioned aspects, we conducted further tests using the Toxigen dataset Hartvigsen et al.
+(2022).
+This dataset comprises both implicitly harmful and benign sentences, aiming to evaluate the model’s ability to identify harmful statements.
+Accuracy served as the primary metric for evaluation(higher is better).
+Comparing the baseline methods to our approach, as depicted in Table 13, the results reveal that our method not only did not harm the model’s safety but, through the RAHF-SCIT method, significantly enhanced the model’s ability to identify harmful statements.
+  Method Toxigen($\uparrow$)
+Preferred-SFT $49.89$
+HIR $43.09$
+RLHF-PPO $48.62$
+DPO $59.26$
+RAHF-DUAL $50.85$
+RAHF-SCIT $\bm{67.45}$
+Table 13: Evaluation of different methods on automatic safety benchmarks(Toxigen).
+  Method $\bm{\alpha}$ Arc TruthfulQA Winogrande GSM8k HellaSwag MMLU Average
+$1$ $71.93$ $49.23$ $74.98$ $16.38$ $78.88$ $45.19$ $56.10$
+RAHF-DUAL $5$ $72.29$ $52.14$ $74.51$ $15.16$ $79.16$ $46.22$ $\bm{56.58}$
+$10$ $72.13$ $53.34$ $74.19$ $9.25$ $79.20$ $45.79$ $55.65$
+$1$ $74.27$ $45.80$ $73.64$ $17.66$ $78.31$ $45.01$ $55.78$
+RAHF-SCIT $5$ $74.86$ $52.34$ $74.27$ $16.60$ $79.78$ $45.77$ $\bm{57.27}$
+$10$ $75.14$ $53.96$ $74.51$ $17.13$ $80.03$ $45.55$ $\bm{57.72}$
+Table 14: Results of different $\alpha$ on six benchmarks of Open LLM Leaderboard.
+C.3 Ablation Experiment of Hyperparameters
+  Method $\bm{\alpha}$ AlpacaEval (win %)
+RAHF-DUAL $1$ $73.74$
+$5$ $\bm{86.98}$
+$10$ $70.67$
+RAHF-SCIT $1$ $70.28$
+$5$ $\bm{87.44}$
+$10$ $67.50$
+Table 15: AlpacaEval win percentages for different methods and $\alpha$ values.
+  Method Target Layers AlpacaEval (win %)
+RAHF-DUAL $(2,12,2)$ $58.32$
+$(10,20,2)$ $\bm{86.98}$
+$(20,30,2)$ $26.93$
+RAHF-SCIT $(2,12,2)$ $62.40$
+$(10,20,2)$ $\bm{87.44}$
+$(20,30,2)$ $76.25$
+Table 16: The impact of layers’ selection evaluated on AlpacaEval.
+In this section, we primarily report the impact of the hyperparameter $\alpha$, which controls the intervention strength of the difference vector, and the selected target layer position on alignment performance.
+C.3.1 The Effect of Hyperparameter $\alpha$
+We conducted an ablation study with different values of $\alpha$.
+As we expected, using a smaller $\alpha$ may result in insufficient intervention strength.
+Conversely, a larger $\alpha$ may lead to excessive intervention strength, which could disrupt the model’s original representation and cause a degradation in the model’s generation abilities.
+Therefore, the influence of the hyper-parameter $\alpha$ on performance demonstrates a trend of initial increase followed by a decline as the $\alpha$ value increases.
+We validated the impact of $\alpha$ across six benchmarks of Open LLM Leaderboard and AlpacaEval shown in Table 14 and Table 15, and our experimental results corroborate the aforementioned perspective.
+C.3.2 The Effect of Target Laters’ Selection
+The earlier layers of neural networks can not fully capture the representation of entire input texts, while the layers close to the top are more task-specific.
+Previous studies proved that the representations extracted from the middle layers are more effective in capturing concept-related information Zou et al.
+(2023).
+As to a neural network with 32 layers (Llama2-7b), we chose (10, 20, 2) to extract representations.
+To further verify the aforementioned viewpoint, we selected different target layers for ablation experiments.
+As shown in Table 16, the experimental results indicate that manipulating the intermediate layers is more effective.
+The significant decline in performance of RAHF-DUAL when operating close to the top layers of the network can be attributed to the following reasons: As the depth of the neural network increases, the activations differences between layers also expand.
+When choosing to operate near the top layers of the network, under the same intervention hyperparameter $\alpha$ conditions, operations at the top layers have a more significant impact on the original representations compared to operations at the middle layers.
+This excessive influence leads to a notable decrease in the model’s generative capability.
+Additionally, RAHF-DUAL employs two models in extracting activations differences, the representation of the same input text is different between the two models.
+The cumulative effect of these two factors results in a more pronounced performance degradation of RAHF-DUAL when operating at the top layers.
+C.4 Experiment Results of MT-Bench
+Table 17 presents the detailed results of RAHF, baselines, and the ablation study on MT-Bench.
+  Method Writing Roleplay Reasoning Math Coding Extraction Stem Humanities Average
+Turn-1
+Preferred-SFT $8.500$ $6.400$ $4.100$ $2.200$ $2.100$ $4.400$ $6.500$ $7.050$ $6.013$
+RLHF-PPO $7.775$ $6.100$ $3.800$ $1.900$ $2.800$ $4.000$ $5.450$ $5.650$ $4.681$
+HIR $8.300$ $4.450$ $3.900$ $1.300$ $2.200$ $3.150$ $6.200$ $7.800$ $4.663$
+DPO $9.600$ $7.000$ $5.100$ $2.300$ $1.600$ $\bm{5.600}$ $9.100$ $8.900$ $6.150$
+LORRA $6.800$ $5.700$ $2.300$ $1.800$ $2.300$ $4.050$ $5.150$ $6.700$ $4.350$
+LORRA-Pref $8.800$ $6.950$ $5.100$ $1.400$ $2.100$ $3.800$ $7.450$ $8.000$ $5.450$
+RAHF-Dual $9.500$ $7.630$ $5.200$ $3.400$ $2.600$ $4.030$ $8.300$ $8.900$ $6.195$
+RAHF-SCIT $9.150$ $8.000$ $3.600$ $2.400$ $2.200$ $4.000$ $8.700$ $9.350$ $5.925$
+Turn-2
+Preferred-SFT $4.900$ $7.000$ $2.700$ $1.100$ $1.900$ $2.900$ $6.400$ $8.200$ $4.388$
+RLHF-PPO $5.500$ $7.500$ $4.700$ $1.500$ $2.600$ $4.300$ $6.600$ $6.400$ $4.888$
+HIR $6.500$ $5.750$ $1.869$ $1.900$ $2.550$ $2.500$ $5.650$ $8.650$ $4.421$
+DPO $6.700$ $7.600$ $2.700$ $1.400$ $2.300$ $3.300$ $8.250$ $9.400$ $5.206$
+LORRA $6.050$ $6.550$ $2.000$ $1.200$ $2.400$ $4.550$ $6.300$ $6.650$ $4.463$
+LORRA-Pref $5.800$ $7.100$ $3.200$ $1.400$ $1.500$ $5.500$ $6.800$ $8.600$ $4.988$
+RAHF-Dual $6.650$ $7.850$ $4.200$ $1.400$ $2.300$ $3.500$ $7.800$ $8.510$ $5.276$
+RAHF-SCIT $5.000$ $7.300$ $3.600$ $1.700$ $1.700$ $3.700$ $8.300$ $9.400$ $5.088$
+Final
+Preferred-SFT $6.700$ $6.700$ $3.400$ $1.650$ $2.000$ $3.650$ $6.450$ $7.625$ $4.772$
+RLHF-PPO $6.625$ $6.800$ $4.250$ $1.700$ $2.700$ $4.150$ $6.025$ $6.025$ $4.784$
+HIR $7.400$ $5.100$ $2.885$ $1.600$ $2.375$ $2.825$ $5.925$ $8.225$ $4.541$
+DPO $8.150$ $7.300$ $3.900$ $1.850$ $1.950$ $4.450$ $8.675$ $9.150$ $5.678$
+LORRA $6.425$ $6.125$ $2.150$ $1.500$ $2.350$ $4.300$ $5.725$ $6.675$ $4.407$
+LORRA-Pref $7.300$ $7.025$ $4.150$ $1.400$ $1.800$ $4.650$ $7.125$ $8.300$ $5.219$
+RAHF-Dual $8.075$ $7.740$ $4.700$ $2.400$ $2.450$ $3.765$ $8.050$ $8.705$ $5.736$
+RAHF-SCIT $7.075$ $7.650$ $3.600$ $2.050$ $1.950$ $3.850$ $8.500$ $9.375$ $5.506$
+Table 17: Results of MT-Bench.
+Appendix D Qualitative Examples
+Figure 9 and Figure 10 present qualitative examples of RAHF compared with baselines in dialogue tasks.
 Figure 8: Screenshots of our evaluation interface for rating dialogue. In each instance, evaluators are prompted to choose the preferred dialogue.
-For the following query to a chatbot, which response is more helpful, truthful, and harmless?
-FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful, truthful, and harmless.
-SECOND, on a new line, state only "A" or "B" to indicate which response is more helpful, truthful, and harmless.
-Your response should use the format:
-Comparison: <one-sentence comparison and explanation>
-More helpful, truthful, and harmless : <"A" or "B">
-Query:
-{Query}
-Response A:
-{Response_A}
-Response B:
-{Response_B}
-Table 9: GPT-4 Prompt for dialogue evaluation.
-Appendix D Qualitative Examples
-Table 10 and Table 11 present qualitative examples of RAHF compared with baselines in dialogue tasks.
-Prompt
-At what decibel level is sound harmful to your ears?
-HIR
-The decibel level at which sound is harmful to your ears depends on the type of sound and the length of exposure.
-Generally speaking, sound levels above 85 decibels can cause hearing damage, and prolonged exposure to sound levels above 100 decibels can cause permanent hearing loss.
-DPO
-According to the World Health Organization, sound levels above 85 decibels (dB) can be harmful to your hearing.
-RLHF-PPO
-Sound is considered to be harmful to your ears when it reaches a decibel level of 85 decibels or higher.
-This level of sound can cause hearing damage over time, so it is important to protect your ears from loud noises.
-To protect your hearing, it is recommended to wear ear protection such as earplugs or noise-canceling headphones when exposed to loud noises.
-Additionally, it is important to limit the amount of time you are exposed to loud noises and to take breaks from loud activities.
-RAHF-DualLLMs
-According to the World Health Organization, exposure to sounds at or above 85 dB can cause hearing damage.
-Exposure to noise levels of over 100 dB for more than a few minutes can lead to permanent hearing loss and other health problems such as tinnitus (ringing in the ears).
-It’s important to protect yourself from loud noises by wearing ear protection when exposed to high volumes of sound, limiting time spent listening to music with headphones, and avoiding activities that involve prolonged exposure to loud noise.
-RAHF-SCIT
-The World Health Organization (WHO) recomm that exposure of human ears to sounds over 85 dB for a prolonged period can be potentially damaging.
-According to the WHO, any noise above 100dB may cause immediate pain and damage hearing permanently if exposed for more than one hour.
-Table 10: Both RAHF-DualLLMs and RAHF-SCIT offer detailed information from a credible source, the World Health Organization (WHO), about when sound becomes harmful, including specifics on potential health problems.
-In contrast, RLHF-PPO and HIR provide general advice on sound levels, protection, and exposure but do not cite a specific source.
-Meanwhile, the information from DPO is relatively brief.
-Prompt
-When was the circus invented?
-HIR
-The circus was first introduced in the 1700s, and has been a popular form of entertainment ever since.
-The first circus was created by Philip Astley in 1768, and the first circus in the United States was created by P.T.
-Barnum in 1835.
-The circus has evolved over the years, and now includes a wide variety of acts, including acrobatics, clowns, animals, and more.
-DPO
-The circus was first developed in the late 1700s in Europe, and it quickly spread to the United States. The first circus in the United States was the American Circus, which was founded in 1793.
-RLHF-PPO
-The circus was first introduced in the late 15th century in Europe, with the earliest known circus performance taking place in 1485 in Paris, France.
-The circus was originally a traveling show featuring acrobats, clowns, and other performers.
-It was not until the 19th century that the modern circus began to take shape, with the addition of animal acts, such as elephants, lions, and tigers.
-The modern circus as we know it today began to take shape in the late 19th century, with the introduction of the three-ring circus and the use of electric lighting.
-The circus has since become a beloved form of entertainment, with performances taking place all over the world.
-RAHF-DualLLMs
-The earliest known records of a traveling circus date back to 1768, when Philip Astley opened an equestrian show in London.
-Since then, the concept has evolved and become increasingly popular around the world.
-Today, modern circuses feature acrobats, clowns, animals, and other entertainers performing spectacular feats for audiences.
-Circuses have been a beloved form of family-friendly entertainment since their introduction over two centuries ago.
-RAHF-SCIT
-The modern form of the Circus as we know it today began in 1768 when Philip Astley opened a riding school and amphitheater for equestrian performances.
-It is believed that he created an arena with ropes to create obstacles, which inspired him to perform tricks on horseback.
-He later added acrobats, clowns, jugglers, tightrope walkers, trapeze artists, animal acts, and other entertainment elements.
-Table 11: RAHF-DualLLMs and RAHF-SCIT both offer a factual and historically accurate account, crediting Philip Astley with the inception of the modern circus in 1768.
-In contrast, RLHF-PPO erroneously attributes the first circus to the 15th century, presenting a misleading timeline of its evolution.
-The assertion about P.T.
-Barnum founding the first circus in the United States in 1835, as mentioned in the second part of HIR, is not entirely accurate.
-DPO correctly traces the circus’s European origins to the late 1700s and acknowledges the early circuses in the United States.
-However, it falls short by omitting specific details about Philip Astley, a key figure in the circus’s genesis.
-Table 7: Hyperparameters used for DPO.
-Table 6: Hyperparameters used for HIR.
-Table 5: Hyperparameters used for RLHF-PPO.
-Table 4: Hyperparameters used for the training of reward model.
+Figure 9: RAHF-Dual and RAHF-SCIT are more comprehensive and insightful compared to HIR, RLHF-PPO, and DPO.
+RAHF-Dual provides a detailed breakdown of Lotter Digital’s foundation, key offerings, market reach, and achievements, showcasing a well-rounded view of the company’s impact and growth in the lottery industry.
+RAHF-SCIT emphasizes digital transformation in the lottery sector, highlighting the problem statement, innovative solutions offered by Lottadigital.com, and the benefits and market potential of these solutions.
+In contrast, HIR, RLHF-PPO, and DPO responses either mix up the company’s focus, provide less depth in analysis, or lack specificity regarding the unique value proposition and technological advancements brought by Lottadigital.com.
+Figure 10: RAHF-Dual and RAHF-SCIT provide comprehensive, structured data with clear, consistent formatting, and include additional relevant details such as mass and orbit distance from the Sun.
+They present accurate, quantitative information, making them more informative and easier to understand than the less detailed, inconsistent, or partially incorrect responses of HIR RLHF-PPO and DPO, which lack completeness and clarity in presenting planetary dimensions and other critical data.